### v1

# Perfect Divisibility and Coloring of Some Bull-Free Graphs

Abstract
A graph $G$ is perfectly divisible if, for each induced subgraph $H$ of $G$, $V(H)$ can be partitioned into $A$ and $B$ such that $H[A]$ is perfect and $\omega(H[B])<\omega(H)$. A bull is a graph consisting of a triangle with two disjoint pendant edge, a fork is a graph obtained from $K_{1,3}$ by subdividing an edge once, and an odd torch is a graph obtained from an odd hole by adding an edge $xy$ such that $x$ is nonadjacent to any vertex on the odd hole and the set of neighbors of $y$ on the odd hole is a stable set.
Chudnovsky and Sivaraman [J. Graph Theory 90 (2019) 54-60] proved that every (odd hole, bull)-free graph and every ($P_{5}$, bull)-free graph are perfectly divisible. Karthick et al. [The Electron. J. of Combin. 29 (2022) P3.19.] proved that every (fork, bull)-free graph is perfectly divisible. Chen and Xu [Discrete Appl. Math. 372 (2025) 298-307.] proved that every ($P_{7},C_{5}$, bull)-free graph is perfectly divisible. Let $H\in${{odd balloon}, $\{P_{8},C_{5}\}\}$. In this paper, we prove that every ($H$, bull)-free graph is perfectly divisible. We also prove that a ($P_{6}$, bull)-free graph is perfectly divisible if and only if it contains no Mycielski-Grötzsch graph as an induced subgraph. As corollaries, these graphs is $\binom{\omega+1}{2}-$colorable. Notice that every odd torch contains an odd hole, a $P_{5}$, and a fork. Therefore, our results generalize the results of these scholars. Moreover, we prove that every ($P_{6}$, bull)-free graph $G$ satisfies $\chi(G)\leq\omega(G)^{7}$.
Key words and phrases: bull-free graphs; perfectly divisibility; chromatic number.
AMS 2000 Subject Classifications: 05C15, 05C75
1 introduction
In this paper, all graphs are finite and simple. As usual, let $P_{k}$ and $C_{k}$ be a path and a cycle on $k$ vertices respectively. We say that a graph $G$ contains a graph $H$ if $H$ is isomorphic to an induced subgraph of $G$. A graph $G$ is $H$-free if it does not contain $H$. Analogously, for a family $\mathcal{H}$ of graphs, we say that $G$ is $\mathcal{H}$-free if $G$ induces no member of $\mathcal{H}$.
A k-coloring of a graph $G=(V,E)$ is a mapping $f$: $V\rightarrow\{1,2,...,k\}$ such that $f(u)\neq f(v)$ whenever $uv\in E$. We say that $G$ is k-colorable if $G$ admits a $k$-coloring. The chromatic number of $G$, denoted by $\chi(G)$, is the smallest positive integer $k$ such that $G$ is $k$-colorable. A clique (resp. stable set) of $G$ is a set of pairwise adjacent (resp. nonadjacent) vertices in $G$. The clique number of $G$, denoted by $\omega(G)$, is the maximum size of a clique in $G$. For a given positive integer $k$, we use the notation $[k]$ to denote the set $\{1,\ldots,k\}$.
[…]
An induced cycle of length $k\geq 4$ is called a hole, and $k$ is the length of the hole. A hole is odd if $k$ is odd, and even otherwise. An antihole is the complement graph of a hole. A graph $G$ is said to be perfect if $\chi(H)=\omega(H)$ for every induced subgraph $H$ of $G$. The famous Strong Perfect Graph Theorem [7] was established by Chudnovsky et al. in 2006:
Theorem 1.1
[7] A graph $G$ is perfect if and only if $G$ is (odd hole, odd antihole)-free.
Let $F$ denote the Mycielski-Grötzsch graph $G_{4}$. A bull is a graph consisting of a triangle with two disjoint pendant edges, an odd torch is a graph obtained from an odd hole by adding an edge $xy$ such that $x$ is nonadjacent to any vertex on the odd hole and the set of neighbors of $y$ on the odd hole is a stable set, a fork is a graph obtained from $K_{1,3}$ by subdividing an edge once, and a E is a graph obtained from an induced $P_{5}=v_{1}v_{2}v_{3}v_{4}v_{5}$ by adding an new vertex exactly adjacent to $v_{3}$ in the $P_{5}$. (See Figure 1)
Figure 1: Illustration of some graphs.
A graph is perfectly divisible if for each induced subgraph $H$ of $G$, $V(H)$ can be partitioned into $A$ and $B$ such that $H[A]$ is perfect and $\omega(H[B])<\omega(H)$. This concept was proposed by Hoáng in [12]. Chudnovsky and Sivaraman [9] proved that every (odd hole, bull)-free graph and every ($P_{5}$, bull)-free graph are perfectly divisible. Karthick et al. [17] proved that every (fork, bull)-free graph is perfectly divisible. Chen and Xu [5] proved that every ($P_{7},C_{5}$, bull)-free graph is perfectly divisible. In this paper, we prove the following theorems.
Theorem 1.2
Let $G$ be a (oddtorch,bull)(odd~torch,bull)-free graph. Then $G$ is perfectly divisible.
Theorem 1.3
Let $G$ be a $(P_{8},C_{5},bull)$-free graph. Then $G$ is perfectly divisible.
It is easy to verify that every odd torch contains an odd hole, a $P_{5}$, a fork, and the graph $E$. Therefore, all the class of (odd hole, bull)-free graphs, ($P_{5}$, bull)-free graphs, (fork, bull)-free, and ($E$, bull)-free graphs are subclasses of the class of (odd balloon, bull)-free graphs. Theorem 1.2 generalizes the results of Chudnovsky and Sivaraman [9] and Karthick et al. [17]; and Theorem 1.3 generalizes the result of Chen and Xu [5].
Recall that we use $F$ to denote the Mycielski-Grötzsch graph. It is certain that $F$ is ($P_{6}$, bull)-free and not perfectly divisible (because $\chi(F)=4$ and $\omega(F)=2$.). In fact, the graph $G$ obtained from $F$ by adding a $K_{n}$ in which each vertex is adjacent to all the vertices in $F$ is ($P_{6}$, bull)-free with clique number $n+2$ and not perfectly divisible. Since every ($P_{5}$, bull)-free graph is perfectly divisible, a natural problem is that under what conditions is ($P_{6}$, bull)-free graph perfectly divisible. In this paper, we prove the following theorem.
Theorem 1.4
Let $G$ be a $(P_{6},\text{bull})$-free graph. Then $G$ is perfectly divisible if and only if $G$ is $F$-free.
[…]
By a simple induction on $\omega(G)$, we have that $\chi(G)\leq\binom{\omega(G)+1}{2}$ for each perfectly divisible graph $G$. According to Theorems 1.2, 1.3 and 1.4, we can directly derive the following corollary.
Corollary 1.1
Let $G$ be a $(H,\text{bull})$-free graph, where $H\in\{\{odd~torch\},\{P_{8},C_{5}\},\{P_{6},F\}\}$ Then $\chi(G)\leq\binom{\omega(G)+1}{2}$.
Corollary 1.2
Let $G$ be a (odd torch, $C_{3}$)-free graph. Then $\chi(G)\leq 3$.
We also consider the binding function of ($P_{6}$, bull)-free graphs, and prove the following theorem.
Theorem 1.5
Let $G$ be a $(P_{6},bull)$-free graph. Then $\chi(G)\leq\omega(G)^{7}$.
We will prove Theorem 1.2 in Section 3, prove Theorems 1.4 and 1.3 in Section 4, and prove Theorem 1.5 in Section 5.
2 Notations and Preliminary Results
For $X\subseteq V(G)$, we use $G[X]$ to denote the subgraph of $G$ induced by $X$. Let $v\in V(G)$, $X\subseteq V(G)$. We use $N_{G}(v)$ to denote the set of vertices adjacent to $v$. Let $d_{G}(v)=|N_{G}(v)|$, $M_{G}(v)=V(G)\setminus(N_{G}(v)\cup\{v\})$, $N_{G}(X)=\{u\in V(G)\setminus X\;|\;u$ has a neighbor in X}X\}, and $M_{G}(X)=V(G)\setminus(X\cup N_{G}(X))$. If it does not cause any confusion, we usually omit the subscript $G$ and simply write $N(v)$, $d(v)$, $M(v)$, $N(X)$ and $M(X)$.
For a subset $A$ of $V(G)$ and a vertex $b\in V(G)\setminus A$, we say that $b$ is complete to $A$ if $b$ is adjacent to every vertex of $A$, and that $b$ is anticomplete to $A$ if $b$ is not adjacent to any vertex of $A$. For two disjoint subsets $A$ and $B$ of $V(G)$, AA is complete to $B$ if every vertex of $A$ is complete to $B$, and $A$ is anticomplete to $B$ if every vertex of $A$ is anticomplete to $B$.
For $A,B\subseteq V(G)$, let $N_{A}(B)=N(B)\cap A$ and $M_{A}(B)=A\setminus(N_{A}(B)\cup B)$. For $u,v\in V(G)$, we simply write $u\sim v$ if $uv\in E(G)$, and write $u\not\sim v$ if $uv\not\in E(G)$.
3 Proof of Theorem 1.2
In this section, we will prove Theorem 1.2. Let $S\subseteq V(G)$ with $1<|S|<|V(G)|$. We say that $S$ is a homogeneous set of $G$ if for any vertex in $V(G)\setminus S$ is either complete to $S$ or anticomplete to $S$. We say that a graph $G$ is locally perfect if for every vertex $v\in V(G)$, $G[N(v)]$ is perfect. The following useful lemmas is important to our proof.
Lemma 3.1
[14] Every minimal nonperfectly divisible graph has no homogeneous set.
Lemma 3.2
[8] If $G$ is a bull-free graph, then either $G$ has a homogeneous set or for every $v\in V(G)$, either $G[N(v)]$ is perfect or $G[M(v)]$ is perfect.
Lemma 3.3
Let $G$ be a connected locally perfect bull-free graph, and let $v\in V(G)$. Then $G[M(v)]$ contains no odd antihole of number of vertices at least 7.
Proof. Suppose to its contrary that $H$ is an odd antihole of $G[M(v)]$ with $V(H)=\{v_{1},v_{2},...,v_{k}\}$, where $k\geq 7$ and $\overline{H}=v_{1}v_{2}\cdots v_{k}v_{1}$. Let $v^{\prime}\in N(v)$. We first prove that
if $y\in N(V(H))$, then $|N(y)\cap V(H)|\geq 2$. (1)
Suppose not, and we may by symmetry assume that $N(y)\cap V(H)=\{v_{1}\}$. But then $\{v_{1},v_{3},v_{5},y,v_{2}\}$ induces a bull, a contradiction. This proves (1).
We next prove that
if $x\in M(V(H))$ and $y\in N(V(H))$ such that $x\sim y$, then $N(y)\cap V(H)$ is a stable set. (2)
Suppose not, and without loss of generality, we may suppose that $v_{1},v_{n}\in N(y)\cap V(H)$ with $v_{1}v_{n}\in E(G)$, where $3\leq n\leq k-2$. We next will show that $y$ is complete to $\{v_{1},v_{2},\cdots,v_{n}\}$. If not, then let $2\leq n^{\prime}\leq n-1$ be the minimum integer such that $y\not\sim v_{n^{\prime}}$. If $n=3$, then $n^{\prime}=2$. To avoid an induced bull on $\{v_{1},v_{3},y,x,v_{4}\}$, we have that $y\sim v_{4}$; but then $\{v_{1},v_{4},y,x,v_{2}\}$ induces a bull, a contradiction. Hence, $n\geq 4$, and thus $v_{n}\sim v_{2}$ and $v_{n-1}\sim v_{1}$.
We can deduce that $n^{\prime}\neq 2$ to avoid an induced bull on $\{v_{1},v_{n},y,x,v_{2}\}$; and $n^{\prime}\neq n-1$ to avoid an induced bull on $\{v_{1},v_{n},y,x,v_{n-1}\}$. We have that $3\leq n^{\prime}\leq n-2$, and so $v_{n^{\prime}}\sim v_{n}$ and $v_{1}\sim v_{n^{\prime}-1}$ by the minimality of $n^{\prime}$. But then $\{v_{n^{\prime}-1},v_{n},y,x,v_{n^{\prime}},\}$ induces a bull, a contradiction. Therefore, $y$ is complete to $\{v_{1},v_{2},\cdots,v_{n}\}$. Similarly, with the same arguments, we can deduce that $y$ is complete to $\{v_{1},v_{k},v_{k-1},\cdots,v_{n}\}$, and this implies that $y$ is complete to $V(H)$, which contradicts to that $G$ is locally perfect. This proves (2).
Since $G$ is connected and $V(H)\subseteq M(v)$, there exist two adjacent vertices $x\in M(V(H))$ and $y\in N(V(H))$. By (1) and (2), we may by symmetry assume that $N(y)\cap V(H)=\{v_{1},v_{2}\}$. But then, $\{v_{k-3},v_{k-1},v_{k},v_{1},y\}$ induces a bull, a contradiction. This completes the proof of Lemma 3.3. $\square$
Lemma 3.4
Let $G$ be a minimal nonperfect divisible bull-free graph. Then $G$ is a connected locally perfect graph, and for every vertex $v\in V(G)$, $G[M(v)]$ is imperfect.
Proof. It is certain that $G$ is connected. Suppose to its contrary that there exists a vertex $v\in V(G)$ such that $G[M(v)]$ is perfect. But then, $G[\{v\}\cup M(v)]$ is perfect and $G[N(v)]$ has clique number less than $\omega(G)$, which contradicts to that $G$ is minimal nonperfectly divisible. Therefore, for every $v\in V(G)$, $G[M(v)]$ is imperfect. Moreover, since $G$ is minimal nonperfectly divisible, $G$ has no homogeneous set by Lemma 3.1. And so, for every $v\in V(G)$, $G[N(v)]$ is perfect by Lemma 3.2. This proves Lemma 3.4. $\square$
By Lemmas 3.3 and 3.4, we have that
Lemma 3.5
Let $G$ be a minimal nonperfectly divisible bull-free graph, and let $v\in V(G)$. Then $G[M(v)]$ contains no odd antihole of number of vertices at least 7.
we now prove Theorem 1.2.
Proof of Theorems 1.2: Suppose the Theorem 1.2 is not true. Let $G$ be a minimal nonperfectly divisible (odd torch, bull)-free graph. By Lemma 3.4, we have that $G$ is connected, and
for every vertex $v\in V(G)$, $G[N(v)]$ is perfect and $G[M(v)]$ is imperfect. (3)
Let $v\in V(G)$. By (3), Lemma 3.5 and Theorem 1.1, $G[M(v)]$ contains an odd hole, say $C=v_{1}v_{2}\cdots v_{k}v_{1}$, where $k\geq 5$ is an odd integer. Since $G$ is connected, there exist two adjacent vertices $x\in M(V(C))$ and $y\in N(V(C))$. Since $G$ is odd torch-free, we have that $N(y)\cap V(C)$ is not a stable set. Hence, we may by symmetry assume that $\{v_{1},v_{2}\}\subseteq N(y)\cap V(C)$.
By (3), we have that $y$ is not complete to $V(C)$ as $G[N(y)]$ is perfect. Let $3\leq n\leq k$ be the minimum integer such that $y\not\sim v_{n}$. But then $\{v_{n-2},v_{n-1},y,x,v_{n}\}$ induces a bull, a contradiction. This completes the proof of Theorem 1.2. $\square$
4 Proof of Theorems 1.3 and 1.4
In this section, we will prove Theorems 1.4 and 1.3. The following useful lemmas is important to our proof.
Lemma 4.1
Let $G$ be a connected locally perfect $(P_{6},bull)$-free graph with $\omega(G)\geq 3$. Let $v\in V(G)$ such that $v$ belongs to a maximum clique of $G$. If $G[M(v)]$ contains a 5-hole, then $G$ has a homogeneous set.
Proof. Let $C=v_{1}v_{2}\cdots v_{5}v_{1}$ be a 5-hole in $G[M(v)]$. From now on, the subscript is modulo 5 in the proof of the Lemma. For $1\leq i\leq 5$, let
$X_{i}$ $=$ $\{u\in N(V(C))~|~N_{C}(u)=\{v_{i}\}\};$
$Y_{i}$ $=$ $\{u\in N(V(C))~|~N_{C}(u)=\{v_{i},v_{i+2}\}\};$
$Z_{i}$ $=$ $\{u\in N(V(C))~|~N_{C}(u)=\{v_{i},v_{i+1},v_{i+2}\}\};$
$W_{i}$ $=$ $\{u\in N(V(C))~|~N_{C}(u)=\{v_{i},v_{i+1},v_{i+2},v_{i+3}\}\};$
Let $X=\bigcup_{i=1}^{5}X_{i}$, $Y=\bigcup_{i=1}^{5}Y_{i}$, $Z=\bigcup_{i=1}^{5}Z_{i}$ and $W=\bigcup_{i=1}^{5}W_{i}$. We next prove the following claim.
Claim 4.1
$N(V(C))=X\cup Y\cup Z\cup W$.
Proof. It suffices to prove that $N(V(C))\subseteq X\cup Y\cup Z\cup W$. Let $u\in N(V(C))$. Without loss of generality, suppose $u\sim v_{1}$. If $N(u)\cap V(C)=\{v_{1}\}$, then $u\in X_{1}$, and we are done. So, suppose $u$ has a neighbor in $\{v_{2},v_{3},v_{4},v_{5}\}$.
If $u$ is anticomplete to $\{v_{2},v_{5}\}$, then $x$ is not complete to $\{v_{3},v_{4}\}$ as otherwise $\{u,v_{3},v_{4},v_{2},v_{5}\}$ induces a bull. Hence, $u$ has exactly one neighbor in $\{v_{3},v_{4}\}$. Now, $u\in Y_{1}$ if $u\sim v_{3}$ and $u\in Y_{4}$ if $u\sim v_{4}$. Therefore, we may assume that $u$ has a neighbor in $\{v_{2},v_{5}\}$, and by symmetry assume $u\sim v_{2}$.
To forbid an induced bull on $\{u,v_{1},v_{2},v_{3},v_{5}\}$, $u$ has a neighbor in $\{v_{3},v_{5}\}$. By symmetry, we may assume $u\sim v_{3}$. If $N(u)\cap V(C)=\{v_{1},v_{2},v_{3}\}$, then $u\in Z_{1}$, and we are done. So, we may suppose $u$ has a neighbor in $\{v_{4},v_{5}\}$. Since $G$ is locally perfect, $u$ is not complete to $V(C)$. Now, $u\in W_{1}$ if $u\sim v_{4}$ and $u\in W_{5}$ if $u\sim v_{5}$. This proves Claim 4.1. $\square$
Consequently, we prove that
for every $u\in N(v)$, $u\in Y$. (4)
Since $G$ is connected, we claim that $u$ has a neighbor in $V(C)$. On the contrary, there exists an induced $P_{3}=x_{1}x_{2}x_{3}$ such that $\{x_{1},x_{2}\}\subseteq M(V(C))$ and $x_{3}\in N(V(C))$. By symmetry, suppose $x_{3}\sim v_{1}$. By Claim 4.1, without loss of generality, we may assume that $x_{3}\in X_{1}\cup Y_{1}\cup Z_{1}\cup W_{1}$. If $x_{3}\in W_{1}$, then $\{x_{3},v_{3},v_{4},x_{2},v_{5}\}$ induces a bull; if $x_{3}\in X_{1}$, then there is an induced $P_{6}=x_{1}x_{2}x_{3}v_{1}v_{2}v_{3}$; otherwise, there is an induced $P_{6}=x_{1}x_{2}x_{3}v_{3}v_{4}v_{5}$. All are contradictions. This implies that $N(u)\cap V(C)\neq\mbox{{\rm\O }}$.
By Claim 4.1, without loss of generality, we assume that $u\in X_{1}\cup Y_{1}\cup Z_{1}\cup W_{1}$. If $u\in X_{1}$, then there is an induced $P_{6}=vuv_{1}v_{2}v_{3}v_{4}$. If $u\in Z_{1}$, then $\{v,u,v_{2},v_{3},v_{4}\}$ induces a bull. If $u\in W_{1}$, then $\{u,v_{3},v_{4},v,v_{5}\}$ induces a bull. So, $u\in Y_{1}$. This proves (4).
We next prove that
if $x,y\in N(v)$ such that $x\sim y$, then $N_{C}(x)=N_{C}(y)$. (5)
Assume for contradiction that $N_{C}(x)\neq N_{C}(y)$. Without loss of generality, suppose $N_{C}(x)=\{v_{1},v_{3}\}$ by (4). If $N_{C}(y)=\{v_{2},v_{4}\}$, then $\{v,x,y,v_{1},v_{4}\}$ induces a bull. If $N_{C}(y)=\{v_{3},v_{5}\}$, then $\{v_{3},x,y,v_{2},v_{5}\}$ induces a bull. By symmetry, in all other cases an induced bull also occurs. Hence, $N_{C}(x)=N_{C}(y)$. This proves (5).
Since $\omega(G)\geq 3$ and $v$ belongs to a maximum clique of $G$, there exist two adjacent vertices in $N(v)$. Let $S$ be a component of $N(v)$ which contains an edge. It is certain that $1<|V(S)|<|V(G)|$. We next prove that $V(S)$ is a homogeneous set of $G$. By (4) and (5), we may by symmetry assume that $N(V(S))\cap V(C)=\{v_{1},v_{3}\}$ as $S$ is connected.
Obviously, we have that $\{v,v_{1},v_{3}\}$ is complete to $V(S)$, and $(N(v)\setminus V(S))\cup\{v_{2},v_{4},v_{5}\}$ is anticomplete to $V(S)$. Suppose $V(S)$ is not a homogeneous set of $G$. There exists a vertex $z$ which has a neighbor and a nonneighbor in $V(S)$. Clearly, $z\notin N(v)\cup\{v\}\cup V(C)$. Since $S$ is connected, there exist an edge $xy\in G[S]$ such that $z\sim x$ and $z\not\sim y$.
To avoid an induced bull on $\{x,y,z,v_{1},v_{2}\}$, we have that $z\sim v_{2}$. Similarly, we have that $z$ is complete to $\{v_{2},v_{4},v_{5}\}$. So $z\in N(V(C))$. By Claim 4.1, $z$ has a neighbor in $\{v_{1},v_{3}\}$. By symmetry, we may assume that $z\sim v_{1}$. But then $\{v_{1},x,z,v,v_{4}\}$ induces a bull, a contradiction. Therefore, $V(S)$ is a homogeneous set of $G$. This completes the proof of Lemma 4.1. $\square$
The proof of following lemma is similar with the proof of Lemma 4.1.
Lemma 4.2
Let $G$ be a connected locally perfect $(P_{8},C_{5},bull)$-free graph with $\omega(G)\geq 3$. Let $v\in V(G)$ such that $v$ belongs to a maximum clique of $G$. If $G[M(v)]$ contains a 7-hole, then $G$ has a homogeneous set.
Proof. Let $C=v_{1}v_{2}\cdots v_{5}v_{6}v_{7}v_{1}$ be a 7-hole in $G[M(v)]$. From now on, the subscript is modulo 7 in the proof of the Lemma. For $1\leq i\leq 7$, let
$X_{i}$ $=$ $\{u\in N(V(C))~|~N_{C}(u)=\{v_{i}\}\};$
$Y_{i}$ $=$ $\{u\in N(V(C))~|~N_{C}(u)=\{v_{i},v_{i+2}\}\};$
$Z_{i}$ $=$ $\{u\in N(V(C))~|~N_{C}(u)=\{v_{i},v_{i+1},v_{i+2}\}\};$
Let $X=\bigcup_{i=1}^{7}X_{i}$, $Y=\bigcup_{i=1}^{7}Y_{i}$ and $Z=\bigcup_{i=1}^{7}Z_{i}$. We next prove the following claim.
Claim 4.2
$N(V(C))=X\cup Y\cup Z$.
Proof. It suffices to prove that $N(V(C))\subseteq X\cup Y\cup Z$. Let $u\in N(V(C))$. Without loss of generality, suppose $u\sim v_{1}$. If $N(u)\cap V(C)=\{v_{1}\}$, then $u\in X_{1}$, and we are done. So, suppose $u$ has a neighbor in $\{v_{2},v_{3},\cdots,v_{7}\}$.
[…]
Without loss of generality, $u\sim v_{2}$. To avoid an induced bull on $\{v_{1},v_{2},u,v_{3},v_{7}\}$, we may by symmetry assume that $u\sim v_{3}$. If $u\sim v_{4}$, then $u\sim v_{5}$ to forbid an induced bull on $\{v_{3},v_{4},u,v_{1},v_{5}\}$; similarly, $u\sim v_{7}$. Since $G$ is locally perfect, $u$ is not complete to $V(C)$, and so $u\not\sim v_{6}$; but then $\{v_{4},v_{5},u,v_{2},v_{6}\}$ induces a bull, a contradiction. So $u\not\sim v_{4}$, and similarly, $u\not\sim v_{7}$.
[…]
Claim 4.3
For every $u\in N(v)$, $u\in Y$. And for $x,y\in N(v)$ with $x\sim y$, $N_{C}(x)=N_{C}(y)$.
Proof. Let $u\in N(v)$. If $N_{C}(u)=\mbox{{\rm\O }}$, since $G$ is connected, there exists an induced $P_{3}=x_{1}x_{2}x_{3}$ such that $\{x_{1},x_{2}\}$ is anticomplete to $V(C)$ and $x_{3}$ has a neighbor in $V(C)$. We may by symmetry assume that $x_{3}\in X_{3}\cup Y_{1}\cup Z_{1}$ by Claim 4.2; but then there exists an induced $P_{8}=x_{!}x_{2}x_{3}v_{3}v_{4}v_{5}v_{6}v_{7}$. So $u\in N(V(C))$. By symmetry and Claim 4.2, we may assume that $u\in X_{3}\cup Y_{1}\cup Z_{1}$. If $u\in X_{1}$, then there is an induced $P_{8}=vuv_{1}v_{2}v_{3}v_{4}v_{5}v_{6}$. If $u\in Z_{1}$, then $\{v_{2},v_{3},u,v,v_{4}\}$ induces a bull. So $u\in Y_{1}\subseteq Y$.
Let $x,y\in N(v)$ such that $x\sim y$. By symmetry, we may suppose that $N_{C}(x)=\{v_{1},v_{3}\}$ and $N_{C}(y)\in\{\{v_{1},v_{3}\},\{v_{2},v_{4}\},\{v_{3},v_{5}\},\{v_{4},v_{6}\}\}$. If $N_{C}(y)=\{v_{2},v_{4}\}$, then $\{v,x,y,v_{1},v_{4}\}$ induces a bull; if $N_{C}(y)=\{v_{3},v_{5}\}$, then $\{v,x,y,v_{1},v_{5}\}$ induces a bull; if $N_{C}(y)=\{v_{4},v_{6}\}$, then $\{v,x,y,v_{1},v_{6}\}$ induces a bull. So, $N_{C}(y)=N_{C}(x)=\{v_{1},v_{3}\}$. This proves Claim 4.3. $\square$
Recall that $v$ belongs to a maximum clique of $G$ and $\omega(G)\geq 3$. Let $S$ be a component of $N(v)$ which contains an edge. It is certain that $1<|V(S)|<|V(G)|$. We next prove that $V(S)$ is a homogeneous set of $G$. By Claim 4.3, we may by symmetry assume that $N(V(S))\cap V(C)=\{v_{1},v_{3}\}$ as $S$ is connected.
It is certain that $\{v,v_{1},v_{3}\}$ is complete to $V(S)$, and $(N(v)\setminus V(S))\cup(V(C)\setminus\{v_{1},v_{3}\})$ is anticomplete to $V(S)$. Suppose that $V(S)$ is not a homogeneous set of $G$. There exists a vertex $z$ which has a neighbor and a nonneighbor in $V(S)$. Clearly, $z\notin N(v)\cup\{v\}\cup V(C)$. Since $S$ is connected, there exist an edge $xy\in G[S]$ such that $z\sim x$ and $z\not\sim y$. To avoid an induced bull on $\{x,y,v_{1},v_{2},z\}$, we have that $z\sim v_{2}$. Similarly, we have that $z$ is complete to $\{v_{2},v_{4},v_{7}\}$. So $z\in N(V(C))$, but $z\notin X\cup Y\cup Z$, which contradicts with Claim 4.2. Therefore, $V(S)$ is a homogeneous set of $G$. This completes the proof of Lemma 4.2. $\square$
We need the following lemmas.
Lemma 4.3
[19] Let $G$ be a $(P_{6},C_{3})$-free graphs. Then $\chi(G)\leq 4$, and $G$ contains Mycielski-Grötzsch graph if $\chi(G)=4$.
Lemma 4.4
[10] Let $G$ be a $(P_{8},C_{5},C_{3})$-free graphs. Then $\chi(G)\leq 3$.
Now, we can prove Theorems 1.3 and 1.4.
Proof of Theorems 1.3 and 1.4: It suffices to prove that every ($P_{8},C_{5}$, bull)-free graph and every ($P_{6}$, bull, $F$)-free graph are perfectly divisible. Suppose not, and let $G$ be a minimal nonperfectly divisible ($H$, bull)-free graph, where $H\in\{P_{6},\{P_{8},C_{5}\}\}$. By Lemmas 3.4, we have that $G$ is a connected locally perfect graph, and for every vertex $v\in V(G)$, $G[M(v)]$ is imperfect. By Lemmas 4.3 and 4.4, we have that $\omega(G)\geq 3$ because $G$ is nonperfectly divisible.
Let $v\in V(G)$ such that $v$ belongs to a maximum clique of $G$. By Lemma 3.5, $G[M(v)]$ is odd antihole-free except $C_{5}$. By Theorem 1.1, we have that $G[M(v)]$ contains an odd hole. Therefore, by Lemmas 4.1 and 4.2, $G$ has a homogeneous set, which contradicts with Lemma 3.1. This completes the proof of Theorems 1.3 and 1.4. $\square$
5 Proof of Theorem 1.5
In this section, we will prove Theorem 1.5. Substituting a vertex $v$ of a graph $G$ by a graph $H$ is an operation that creates a graph obtained from the disjoint union of $H$ and GvG-v by adding an edge between every vertex of $H$ and every neighbor of $v$ in $G$. Let $\cal C$} be a class of graphs. We denote by ${\cal C}_{s}$ the closure of $\cal C$} under taking substitution.
We say a bull-free graph $G$ is basic if neither $G$ nor $\overline{G}$ contains an odd hole with both a vertex which is complete to the odd hole and a vertex which is anticomplete to the odd hole. Chudnovsky and Safra [8] proved the following two lemmas.
Lemma 5.1
((1.4 of [8] )) Every bull-free graph can be obtained via substitution from basic bull-free graphs.
Lemma 5.2
((4.3 of [8] )) Let $G$ be a basic bull-free graph. Then for every vertex $v\in V(G)$, either $G[N(v)]$ is perfect or $G[M(v)]$ is perfect.
Lemma 5.3
((6.1 of [2] )) Let $\cal C$} be a class of graphs. If $\cal C$} is hereditary and Polynomially $\chi$-bounded with $\chi\leq\omega^{k}$, then ${\cal C}_{s}$ is polynomially $\chi$-bounded with $\chi\leq\omega^{2k+3}$.
Lemma 5.4
[18] The class of perfect graphs is closed under taking substitution.
Lemma 5.5
Let $G$ be a locally perfect $(P_{6},bull)$-free graph. Then there exists a partition $(X,Y)$ of $V(G)$ such that $G[X]$ and $G[Y]$ are perfect.
Proof. We prove it by induction on $|V(G)|$. It is certain that it holds when $|V(G)|$ is small. Now we proceed with the inductive proof. If $\omega(G)\leq 2$, by Lemma 4.3, then $\chi(G)\leq 4$ and we are done. Therefore, by inductive hypothesis, we may assume that
$G$ is connected and $\omega(G)\geq 3$. (6)
Let $v\in V(G)$ such that $v$ belongs to a maximum clique of $G$. If $G[M(v)]$ is perfect, then $(N(v),M(v)\cup\{v\})$ is a partition of $V(G)$ which satisfies the Lemma since $G$ is locally perfect. Therefore, we may assume that $G[M(v)]$ is imperfect. By Lemma 3.3, $G[M(v)]$ contains no odd antihole of number of vertices at least 7. So, by Theorem 1.1, $G[M(v)]$ contains a 5-hole as $G$ is $P_{6}$-free. By Lemma 4.1, we have that $G$ has a homogeneous set, say $S$. Recall that $1<|S|<|V(G)|$.
Since $G$ is connected, we have that there exists a vertex which is complete to $S$, and so $G[S]$ is perfect because $G$ is locally perfect. Since $S\neq\mbox{{\rm\O }}$, we may assume that $x\in S$. Let $G^{\prime}=G[V(G)\setminus(S\setminus\{x\})]$. It is easy to verify that $G^{\prime}$ is a locally perfect $(P_{6}$, bull)-free graph and $|V(G^{\prime})|<|V(G)|$. Therefore, by inductive hypothesis, there exists a partition $(X^{\prime},Y^{\prime})$ of $V(G^{\prime})$ such that $G[X^{\prime}]$ and $G[Y^{\prime}]$ is perfect. Since $x\in V(G^{\prime})$, without loss of generality, we may assume that $x\in X^{\prime}$. Let $X=X^{\prime}\cup S$ and $Y=Y^{\prime}$. Clearly, $(X,Y)$ is a partition of $V(G)$. Since $S$ is a homogeneous set of $G$, we have that $S$ is a homogeneous set of $G[X]$, and thus $G[X]$ is a graph obtained from $G[X^{\prime}]$ by substituting the vertex $x$ by the graph $G[S]$. By Lemma 5.4, $G[X]$ is perfect. Therefore, the partition $(X,Y)$ satisfies the Lemma. This completes the proof of Lemma 5.5. $\square$
Lemma 5.6
Let $G$ be a basic $(P_{6},bull)$-free graph. Then $\chi(G)\leq\omega(G)^{2}$.
Proof. We prove it by induction on $|V(G)|$. It is certain that it holds when $|V(G)|$ is small. Now we proceed with the inductive proof. We may assume that $\omega(G)\geq 2$ and $G$ is connected by inductive hypothesis.
If $G$ is locally perfectly, then $\chi(G)\leq 2\omega(G)\leq\omega(G)^{2}$ by Lemma 5.5, and so we are done. Therefore, we may assume that $G$ is not locally perfect. Now, by Lemma 5.1, there exists a vertex $v$ such that $G[M(v)]$ is perfect. By inductive hypothesis, we have that $\chi(G[N(v)])\leq(\omega(G)-1)^{2}$. Therefore, we have that
$\chi(G)\leq\chi(G[N(v)])+\chi(G[\{v\}\cup M(v)])\leq(\omega(G)-1)^{2}+\omega(G)\leq\omega(G)^{2}.$
This completes the proof of Lemma 5.6. $\square$
Now, we can prove Theorem 1.5.
Proof of Theorem 1.5: Let $\cal G$} be the class of ($P_{6}$, bull)-free graphs, and $\cal C$} be the class of basic ($P_{6}$, bull)-free graphs. It is easy to verify that $\cal C$} is hereditary. By Lemmas 5.3 and 5.6, it suffices to show that ${\cal G}\subseteq{\cal C}_{s}$.
Let $G\in{\cal G}$. Since $G$ is bull-free, by Lemma 5.1, we have that $G$ can be obtained from a basic bull-free graph $G_{1}$ by substituting a vertex by a basic bull-free graph $G_{2}$. Since $G$ is $P_{6}$-free and $G_{1},G_{2}$ are induced subgraph of $G$, we have that both $G_{1}$ and $G_{2}$ are $P_{6}$-free; this implies that $G_{1},G_{2}\in{\cal C}$. Therefore, $G\in{\cal C}_{s}$, and so ${\cal G}\subseteq{\cal C}_{s}$.
This completes the proof of Theorem 1.5. $\square$
Remark
[…]
More generally, let ${\cal G}_{1}$ be a hereditary class of graphs in which every trianlge-free graph is $k$-colorable, let ${\cal G}_{2}$ be the class of bull-free graphs, and let ${\cal G}={\cal G}_{1}\cap{\cal G}_{2}$. In [6], Chudnovsky et al. proved that $\cal G$} is polynomially $\chi$-bounded. Here, what are we more concerned about is that is it true that every graph in $\cal G$} is perfectly divisible if and only if $k=3$.
Declarations
  • Research of the first author was supported by Postgraduate Research and Practice Innovation Program of Jiangsu Province KYCX25_1926. And research of the second author was supported by the Scientific Research Foundation of Nanjing Institute of Technology, China (No. YKJ202448).
  • Conflict of interest The authors declare no conflict of interest.
  • Data availibility statement This manuscript has no associated data.

### v3

@@
 A graph $G$ is perfectly divisible if, for each induced subgraph $H$ of $G$, $V(H)$ can be partitioned into $A$ and $B$ such that $H[A]$ is perfect and $\omega(H[B])<\omega(H)$.
-A bull is a graph consisting of a triangle with two disjoint pendant edge, a fork is a graph obtained from $K_{1,3}$ by subdividing an edge once, and an odd torch is a graph obtained from an odd hole by adding an edge $xy$ such that $x$ is nonadjacent to any vertex on the odd hole and the set of neighbors of $y$ on the odd hole is a stable set.
+A bull is a graph consisting of a triangle with two disjoint pendant edges, a fork is a graph obtained from $K_{1,3}$ by subdividing an edge once, and an odd torch is a graph obtained from an odd hole by adding an edge $xy$ such that $x$ is non-adjacent to any vertex on the odd hole and the set of neighbors of $y$ on the odd hole is a stable set.
 Chudnovsky and Sivaraman [J.
@@
 372 (2025) 298-307.] proved that every ($P_{7},C_{5}$, bull)-free graph is perfectly divisible.
-Let $H\in${{odd balloon}, $\{P_{8},C_{5}\}\}$.
+Let $H\in${{odd torch}, $\{P_{8},C_{5}\}\}$.
 In this paper, we prove that every ($H$, bull)-free graph is perfectly divisible.
 We also prove that a ($P_{6}$, bull)-free graph is perfectly divisible if and only if it contains no Mycielski-Grötzsch graph as an induced subgraph.
-As corollaries, these graphs is $\binom{\omega+1}{2}-$colorable.
-Notice that every odd torch contains an odd hole, a $P_{5}$, and a fork.
-Therefore, our results generalize the results of these scholars.
+As corollaries, these graphs are $\binom{\omega+1}{2}-$colorable.
+Notice that every odd torch contains an odd hole, an induced $P_{5}$, and an induced fork.
+Therefore, our results generalize their findings.
 Moreover, we prove that every ($P_{6}$, bull)-free graph $G$ satisfies $\chi(G)\leq\omega(G)^{7}$.
@@
 In this paper, all graphs are finite and simple.
+We follow [1] for undefined notations and terminology.
 As usual, let $P_{k}$ and $C_{k}$ be a path and a cycle on $k$ vertices respectively.
@@
 Analogously, for a family $\mathcal{H}$ of graphs, we say that $G$ is $\mathcal{H}$-free if $G$ induces no member of $\mathcal{H}$.
-A k-coloring of a graph $G=(V,E)$ is a mapping $f$: $V\rightarrow\{1,2,...,k\}$ such that $f(u)\neq f(v)$ whenever $uv\in E$.
+For a given positive integer $k$, we use the notation $[k]$ to denote the set $\{1,\ldots,k\}$.
+A k-coloring of a graph $G=(V,E)$ is a mapping $f$: $V\rightarrow[k]$ such that $f(u)\neq f(v)$ whenever $uv\in E$.
 We say that $G$ is k-colorable if $G$ admits a $k$-coloring.
@@
 stable set) of $G$ is a set of pairwise adjacent (resp.
-nonadjacent) vertices in $G$.
+non-adjacent) vertices in $G$.
 The clique number of $G$, denoted by $\omega(G)$, is the maximum size of a clique in $G$.
-For a given positive integer $k$, we use the notation $[k]$ to denote the set $\{1,\ldots,k\}$.
 […]
@@
 Theorem 1.1
-[7] A graph $G$ is perfect if and only if $G$ is (odd hole, odd antihole)-free.
+[7] A graph $G$ is perfect if and only if $G$ is ((odd hole, odd antihole))-free.
 Let $F$ denote the Mycielski-Grötzsch graph $G_{4}$.
-A bull is a graph consisting of a triangle with two disjoint pendant edges, an odd torch is a graph obtained from an odd hole by adding an edge $xy$ such that $x$ is nonadjacent to any vertex on the odd hole and the set of neighbors of $y$ on the odd hole is a stable set, a fork is a graph obtained from $K_{1,3}$ by subdividing an edge once, and a E is a graph obtained from an induced $P_{5}=v_{1}v_{2}v_{3}v_{4}v_{5}$ by adding an new vertex exactly adjacent to $v_{3}$ in the $P_{5}$.
+A bull is a graph consisting of a triangle with two disjoint pendant edges, an odd torch is a graph obtained from an odd hole by adding an edge $xy$ such that $x$ is non-adjacent to any vertex on the odd hole and the set of neighbors of $y$ on the odd hole is a stable set, a fork is a graph obtained from $K_{1,3}$ by subdividing an edge once, and a E is a graph obtained from an induced $P_{5}=v_{1}v_{2}v_{3}v_{4}v_{5}$ by adding an new vertex exactly adjacent to $v_{3}$ in the $P_{5}$.
 (See Figure 1)
@@
 Karthick et al.
-[17] proved that every (fork, bull)-free graph is perfectly divisible.
+[15] proved that every (fork, bull)-free graph is perfectly divisible, and this class of graphs was generalized by Wu and Xu [19].
 Chen and Xu [5] proved that every ($P_{7},C_{5}$, bull)-free graph is perfectly divisible.
@@
 Let $G$ be a $(P_{8},C_{5},bull)$-free graph. Then $G$ is perfectly divisible.
-It is easy to verify that every odd torch contains an odd hole, a $P_{5}$, a fork, and the graph $E$.
-Therefore, all the class of (odd hole, bull)-free graphs, ($P_{5}$, bull)-free graphs, (fork, bull)-free, and ($E$, bull)-free graphs are subclasses of the class of (odd balloon, bull)-free graphs.
+It is easy to verify that every odd torch contains an odd hole, an induced $P_{5}$, an induced fork, and the graph $E$.
+Therefore, all the class of (odd hole, bull)-free graphs, ($P_{5}$, bull)-free graphs, (fork, bull)-free, and ($E$, bull)-free graphs are subclasses of the class of (odd torch, bull)-free graphs.
 Theorem 1.2 generalizes the results of Chudnovsky and Sivaraman [9] and Karthick et al.
@@
 In fact, the graph $G$ obtained from $F$ by adding a $K_{n}$ in which each vertex is adjacent to all the vertices in $F$ is ($P_{6}$, bull)-free with clique number $n+2$ and not perfectly divisible.
-Since every ($P_{5}$, bull)-free graph is perfectly divisible, a natural problem is that under what conditions is ($P_{6}$, bull)-free graph perfectly divisible.
+Since every ($P_{5}$, bull)-free graph is perfectly divisible, a natural problem is that under what conditions is ($P_{6}$, bull)-free graph perfectly divisible ?
 In this paper, we prove the following theorem.
@@
 Corollary 1.1
-Let $G$ be a $(H,\text{bull})$-free graph, where $H\in\{\{odd~torch\},\{P_{8},C_{5}\},\{P_{6},F\}\}$ Then $\chi(G)\leq\binom{\omega(G)+1}{2}$.
+Let $G$ be an $(H,\text{bull})$-free graph, where $H\in\{\{odd~torch\},\{P_{8},C_{5}\},\{P_{6},F\}\}$ Then $\chi(G)\leq\binom{\omega(G)+1}{2}$.
 Corollary 1.2
-Let $G$ be a (odd torch, $C_{3}$)-free graph. Then $\chi(G)\leq 3$.
+Let $G$ be an ((odd torch, $C_{3}$))-free graph. Then $\chi(G)\leq 3$.
 We also consider the binding function of ($P_{6}$, bull)-free graphs, and prove the following theorem.
@@
 Let $G$ be a $(P_{6},bull)$-free graph. Then $\chi(G)\leq\omega(G)^{7}$.
+Let $G$ be a graph and let $w$ be a weight function on $V(G)$.
+We use $\omega_{w}(G)$ to denote the maximum weight of a clique in $G$.
+A graph $G$ is perfectly weight divisible [9, 12] if for every positive integer weight function $w$ on $V(G)$, and for every induced subgraph $H$ of $G$, there is a partition of $V(H)$ into $P$, $W$ such that $H[P]$ is perfect and $\omega_{w}(H[W])<\omega_{w}(H)$.
+It is certain that if a graph $G$ is perfectly weight divisible, then $G$ is perfectly divisible.
+If $G$ is not perfectly weight divisible but each of its proper induced subgraph is, then we say $G$ is minimally non-perfectly weight divisible (MNWD for short).
+As mentioned above, the structure and coloring problem of bull-free graphs has garnered significant attention, with new methods and research content [4, 6, 11, 13, 18] being continuously proposed.
+This paper utilizes the concept of perfectly weight divisibility to further analyze the structure of bull-free graphs and bull-free MNWD graphs, thereby deriving our main results.
 We will prove Theorem 1.2 in Section 3, prove Theorems 1.4 and 1.3 in Section 4, and prove Theorem 1.5 in Section 5.
@@
 3 Proof of Theorem 1.2
-In this section, we will prove Theorem 1.2.
 Let $S\subseteq V(G)$ with $1<|S|<|V(G)|$.
-We say that $S$ is a homogeneous set of $G$ if for any vertex in $V(G)\setminus S$ is either complete to $S$ or anticomplete to $S$.
+We say that $S$ is a homogeneous set of $G$ if every vertex in $V(G)\setminus S$ is either complete to $S$ or anticomplete to $S$.
 We say that a graph $G$ is locally perfect if for every vertex $v\in V(G)$, $G[N(v)]$ is perfect.
-The following useful lemmas is important to our proof.
+In this section, we will prove Theorem 1.2. The following useful lemmas is important to our proof.
 Lemma 3.1
-[14] Every minimal nonperfectly divisible graph has no homogeneous set.
+[9] Every MNWD graph has no homogeneous set.
 Lemma 3.2
@@
 Suppose not, and without loss of generality, we may suppose that $v_{1},v_{n}\in N(y)\cap V(H)$ with $v_{1}v_{n}\in E(G)$, where $3\leq n\leq k-2$.
-We next will show that $y$ is complete to $\{v_{1},v_{2},\cdots,v_{n}\}$.
+We next show that $y$ is complete to $\{v_{1},v_{2},\cdots,v_{n}\}$.
 If not, then let $2\leq n^{\prime}\leq n-1$ be the minimum integer such that $y\not\sim v_{n^{\prime}}$.
@@
 Lemma 3.4
-Let $G$ be a minimal nonperfect divisible bull-free graph. Then $G$ is a connected locally perfect graph, and for every vertex $v\in V(G)$, $G[M(v)]$ is imperfect.
+Let $G$ be a bull-free MNWD graph. Then $G$ is a connected locally perfect graph, and for every vertex $v\in V(G)$, $G[M(v)]$ is imperfect.
 Proof.
 It is certain that $G$ is connected.
-Suppose to its contrary that there exists a vertex $v\in V(G)$ such that $G[M(v)]$ is perfect.
-But then, $G[\{v\}\cup M(v)]$ is perfect and $G[N(v)]$ has clique number less than $\omega(G)$, which contradicts to that $G$ is minimal nonperfectly divisible.
+Since $G$ is MNWD, we have that for every vertex $v\in V(G)$, $G[M(v)]$ cannot be perfect as otherwise, $(\{v\}\cup M(v),N(v))$ is a partition of $V(G)$ such that for all positive integer weight function on $V(G)$, the $G[\{v\}\cup M(v)]$ is a perfect graph and $G[N(v)]$ has a smaller maximum weight of a clique than $G$, which contradicts with the minimality of $G$.
 Therefore, for every $v\in V(G)$, $G[M(v)]$ is imperfect.
-Moreover, since $G$ is minimal nonperfectly divisible, $G$ has no homogeneous set by Lemma 3.1.
-And so, for every $v\in V(G)$, $G[N(v)]$ is perfect by Lemma 3.2.
+Moreover, since $G$ is MNWD, $G$ has no homogeneous set by Lemma 3.1.
+Therefore, for every $v\in V(G)$, $G[N(v)]$ is perfect by Lemma 3.2.
 This proves Lemma 3.4.
@@
 Lemma 3.5
-Let $G$ be a minimal nonperfectly divisible bull-free graph, and let $v\in V(G)$. Then $G[M(v)]$ contains no odd antihole of number of vertices at least 7.
+Let $G$ be a bull-free MNWD graph, and let $v\in V(G)$. Then $G[M(v)]$ contains no odd antihole of number of vertices at least 7.
 We now prove Theorem 1.2.
-Proof of Theorems 1.2: Suppose the Theorem 1.2 is not true. Let $G$ be a minimal nonperfectly divisible (odd torch, bull)-free graph. By Lemma 3.4, we have that $G$ is connected, and
+Proof of Theorems 1.2: It suffices to show that each (odd torch, bull)-free graph is perfectly weight divisible.
+Suppose it is not true, and let $G$ be a (odd torch, bull)-free MNWD graph.
+By Lemma 3.4, we have that $G$ is connected, and
 for every vertex $v\in V(G)$, $G[N(v)]$ is perfect and $G[M(v)]$ is imperfect. (3)
@@
 But then $\{v_{n-2},v_{n-1},y,x,v_{n}\}$ induces a bull, a contradiction.
+Therefore, each (odd torch, bull)-free graph is perfectly weight divisible.
 This completes the proof of Theorem 1.2.
@@
 4 Proof of Theorems 1.3 and 1.4
-In this section, we will prove Theorems 1.4 and 1.3. The following useful lemmas is important to our proof.
+In this section, we will prove Theorems 1.3 and 1.4. The following useful lemmas are important to our proof.
+Lemma 4.1
+Let $G$ be a graph with $\chi(G)\leq 3$. Then $G$ is perfectly weight divisible.
+Proof.
+Let $w$ be a positive integer weight function on $V(G)$, and let $G^{\prime}$ be an induced subgraph of $G$.
+Since $\chi(G^{\prime})\leq 3$, there is a partition $(A,B)$ of $V(G^{\prime})$ such that $G^{\prime}[A]$ is bipartite and $B$ is a stable set of $G^{\prime}$.
+Let $X\subseteq B$ such that each vertex in $X$ is a isolated vertex of $G^{\prime}$.
+Then $(A\cup X,B\setminus X)$ is a partition of $G^{\prime}$ such that $G[A\cup X]$ is perfect, and $\omega_{w}(G[B\setminus X])<\omega_{w}(G^{\prime})$.
+This proves the lemma.
+$\square$
 Lemma 4.2
@@
 Now, $u\in Y_{1}$ if $u\sim v_{3}$ and $u\in Y_{4}$ if $u\sim v_{4}$.
-Therefore, we may assume that $u$ has a neighbor in $\{v_{2},v_{5}\}$, and by symmetry assume $u\sim v_{2}$.
+Therefore, we may assume that $u$ has a neighbor in $\{v_{2},v_{5}\}$, and by symmetry assume that $u\sim v_{2}$.
 To forbid an induced bull on $\{u,v_{1},v_{2},v_{3},v_{5}\}$, $u$ has a neighbor in $\{v_{3},v_{5}\}$.
-By symmetry, we may assume $u\sim v_{3}$.
+By symmetry, we may assume that $u\sim v_{3}$.
 If $N(u)\cap V(C)=\{v_{1},v_{2},v_{3}\}$, then $u\in Z_{1}$, and we are done.
-So, we may suppose $u$ has a neighbor in $\{v_{4},v_{5}\}$.
-Since $G$ is locally perfect, $u$ is not complete to $V(C)$.
+So, we may suppose that $u$ has a neighbor in $\{v_{4},v_{5}\}$.
+Since $G$ is locally perfect, we have that $u$ is not complete to $V(C)$.
 Now, $u\in W_{1}$ if $u\sim v_{4}$ and $u\in W_{5}$ if $u\sim v_{5}$.
@@
 Obviously, we have that $\{v,v_{1},v_{3}\}$ is complete to $V(S)$, and $(N(v)\setminus V(S))\cup\{v_{2},v_{4},v_{5}\}$ is anticomplete to $V(S)$.
-Suppose $V(S)$ is not a homogeneous set of $G$.
-There exists a vertex $z$ which has a neighbor and a nonneighbor in $V(S)$.
+Suppose that $V(S)$ is not a homogeneous set of $G$.
+There exists a vertex $z$ which has a neighbor and a non-neighbor in $V(S)$.
 Clearly, $z\notin N(v)\cup\{v\}\cup V(C)$.
@@
 $\square$
-The proof of following lemma is similar with the proof of Lemma 4.1.
+The proof of the following lemma is similar to that of Lemma 4.2.
 Lemma 4.3
@@
 If $G[M(v)]$ contains a 7-hole, then $G$ has a homogeneous set.
-Proof. Let $C=v_{1}v_{2}\cdots v_{5}v_{6}v_{7}v_{1}$ be a 7-hole in $G[M(v)]$. From now on, the subscript is modulo 7 in the proof of the Lemma. For $1\leq i\leq 7$, let
+Proof. Let $C=v_{1}v_{2}\cdots v_{5}v_{6}v_{7}v_{1}$ be a 7-hole in $G[M(v)]$. For the remainder of the proof, all subscripts are taken modulo 7. For $1\leq i\leq 7$, let
 $X_{i}$ $=$ $\{u\in N(V(C))~|~N_{C}(u)=\{v_{i}\}\};$
@@
 So, suppose $u$ has a neighbor in $\{v_{2},v_{3},\cdots,v_{7}\}$.
-[…]
+Suppose that $u$ is anticomplete to $\{v_{2},v_{7}\}$.
+Then $u$ has a neighbor in $\{v_{3},v_{4},v_{5},v_{6}\}$.
+Assume that $u$ is anticomplete to $\{v_{3},v_{6}\}$.
+If $u\sim v_{4}$, then there is a 5-hole $uv_{1}v_{2}v_{3}v_{4}u$, and we are done by Lemma 4.2.
+So, $u\not\sim v_{4}$; similarly, $u\not\sim v_{5}$, a contradiction.
+Therefore, $u$ is not anticomplete to $\{v_{3},v_{6}\}$ and we may by symmetry assume that $u\sim v_{3}$.
+If $u\sim v_{6}$, then $u$ has a neighbor in $\{v_{4},v_{5}\}$ to avoid a 5-hole $uv_{3}v_{4}v_{5}v_{6}u$; and we may by symmetry assume that $u\sim v_{4}$, but now $\{v_{3},v_{4},u,v_{2},v_{6}\}$ induces a bull.
+So, $u\not\sim v_{6}$.
+To avoid a 5-hole $uv_{5}v_{6}v_{7}v_{1}u$, we have that $u\not\sim v_{5}$.
+Moreover, to avoid an induced bull on $\{v_{3},v_{4},u,v_{2},v_{5}\}$, we have that $u\not\sim v_{4}$.
+Now, $N(u)\cap V(C)=\{v_{1},v_{3}\}$, then $u\in Y_{1}$ and we are done.
+So, $u$ is not anticomplete to $\{v_{2},v_{7}\}$.
 Without loss of generality, $u\sim v_{2}$.
@@
 So $u\not\sim v_{4}$, and similarly, $u\not\sim v_{7}$.
-[…]
+If $u$ is complete to $\{v_{5},v_{6}\}$, then $\{v_{5},v_{6},u,v_{4},v_{7}\}$ induces a bull.
+So $u$ has at most one neighbor in $\{v_{5},v_{6}\}$.
+If $u\sim v_{5}$, then there is a 5-hole $uv_{5}v_{6}v_{7}v_{1}u$.
+Hence, $u$ is anticomplete to $\{v_{5},v_{6}\}$ by symmetry.
+And thus $N(u)\cap V(C)=\{v_{1},v_{2},v_{3}\}$; this implies $u\in Z_{1}$.
+This proves Claim 4.2.
+$\square$
 Claim 4.3
-For every $u\in N(v)$, $u\in Y$. And for $x,y\in N(v)$ with $x\sim y$, $N_{C}(x)=N_{C}(y)$.
+For every $u\in N(v)$, we have that $u\in Y$. And for $x,y\in N(v)$ with $x\sim y$, $N_{C}(x)=N_{C}(y)$.
 Proof.
@@
 Suppose that $V(S)$ is not a homogeneous set of $G$.
-There exists a vertex $z$ which has a neighbor and a nonneighbor in $V(S)$.
+There exists a vertex $z$ which has a neighbor and a non-neighbor in $V(S)$.
 Clearly, $z\notin N(v)\cup\{v\}\cup V(C)$.
@@
 Now, we can prove Theorems 1.3 and 1.4.
-Proof of Theorems 1.3 and 1.4: It suffices to prove that every ($P_{8},C_{5}$, bull)-free graph and every ($P_{6}$, bull, $F$)-free graph are perfectly divisible.
-Suppose not, and let $G$ be a minimal nonperfectly divisible ($H$, bull)-free graph, where $H\in\{P_{6},\{P_{8},C_{5}\}\}$.
-By Lemmas 3.4, we have that $G$ is a connected locally perfect graph, and for every vertex $v\in V(G)$, $G[M(v)]$ is imperfect.
-By Lemmas 4.3 and 4.4, we have that $\omega(G)\geq 3$ because $G$ is nonperfectly divisible.
+Proof of Theorems 1.3 and 1.4: Since the graph $F$ is non-perfectly divisible, it suffices to prove that every ($P_{8},C_{5}$, bull)-free graph and every ($P_{6}$, bull, $F$)-free graph are perfectly divisible.
+We will actually prove the following stronger conclusion: every ($P_{8},C_{5}$, bull)-free graph and every ($P_{6}$, bull, $F$)-free graph are perfectly weight divisible.
+Suppose not, and let $G$ be a ($H$, bull)-free MNWD graph, where $H\in\{P_{6},\{P_{8},C_{5}\}\}$.
+By Lemma 3.4, we have that $G$ is a connected locally perfect graph, and for every vertex $v\in V(G)$, $G[M(v)]$ is imperfect.
+By Lemmas 4.14.4 and 4.5, we have that $\omega(G)\geq 3$ because $G$ is non-perfectly weight divisible.
 Let $v\in V(G)$ such that $v$ belongs to a maximum clique of $G$.
-By Lemma 3.5, $G[M(v)]$ is odd antihole-free except $C_{5}$.
+By Lemma 3.5, $G[M(v)]$ contains no odd antihole of number of vertices at least 7.
 By Theorem 1.1, we have that $G[M(v)]$ contains an odd hole.
-Therefore, by Lemmas 4.1 and 4.2, $G$ has a homogeneous set, which contradicts with Lemma 3.1.
+Therefore, by Lemmas 4.2 and 4.3, $G$ has a homogeneous set, which contradicts with Lemma 3.1.
 This completes the proof of Theorems 1.3 and 1.4.
@@
 In this section, we will prove Theorem 1.5.
+A class of graphs is hereditary if it is closed under taking induced subgraphs.
 Substituting a vertex $v$ of a graph $G$ by a graph $H$ is an operation that creates a graph obtained from the disjoint union of $H$ and GvG-v by adding an edge between every vertex of $H$ and every neighbor of $v$ in $G$.
@@
 Since $G$ is bull-free, by Lemma 5.1, we have that $G$ can be obtained from a basic bull-free graph $G_{1}$ by substituting a vertex by a basic bull-free graph $G_{2}$.
-Since $G$ is $P_{6}$-free and $G_{1},G_{2}$ are induced subgraph of $G$, we have that both $G_{1}$ and $G_{2}$ are $P_{6}$-free; this implies that $G_{1},G_{2}\in{\cal C}$.
+Since $G$ is $P_{6}$-free and $G_{1},G_{2}$ are induced subgraphs of $G$, we have that both $G_{1}$ and $G_{2}$ are $P_{6}$-free, which implies that $G_{1},G_{2}\in{\cal C}$.
 Therefore, $G\in{\cal C}_{s}$, and so ${\cal G}\subseteq{\cal C}_{s}$.
@@
 proved that $\cal G$} is polynomially $\chi$-bounded.
-Here, what are we more concerned about is that is it true that every graph in $\cal G$} is perfectly divisible if and only if $k=3$.
+Here, what are we more concerned about is that is it true that every graph in $\cal G$} is perfectly divisible if and only if $k=3$ ?
 Declarations
-  • Research of the first author was supported by Postgraduate Research and Practice Innovation Program of Jiangsu Province KYCX25_1926.
-And research of the second author was supported by the Scientific Research Foundation of Nanjing Institute of Technology, China (No.
+  • Ran Chen is supported by Postgraduate Research and Practice Innovation Program of Jiangsu Province KYCX25_1926.
+Di Wu is supported by the Scientific Research Foundation of Nanjing Institute of Technology, China (No.
 YKJ202448).

### v5

@@
-Perfect Divisibility and Coloring of Some Bull-Free Graphs
+On Perfect Divisibility of Bull-Free Graphs Without Long Paths
 Abstract
 A graph $G$ is perfectly divisible if, for each induced subgraph $H$ of $G$, $V(H)$ can be partitioned into $A$ and $B$ such that $H[A]$ is perfect and $\omega(H[B])<\omega(H)$.
-A bull is a graph consisting of a triangle with two disjoint pendant edges, a fork is a graph obtained from $K_{1,3}$ by subdividing an edge once, and an odd torch is a graph obtained from an odd hole by adding an edge $xy$ such that $x$ is non-adjacent to any vertex on the odd hole and the set of neighbors of $y$ on the odd hole is a stable set.
+A bull is a graph consisting of a triangle with two disjoint pendant edges.
+Let $H$ be an odd hole or an odd antihole.
+An odd net is the graph obtained from $H$ by adding a new vertex adjacent to all vertices of $H$.
+An odd net+ is the graph obtained from an odd net by adding an additional new vertex that is non-adjacent to any vertex of $H$.
+Let $F$ denote the Grötzsch graph.
 Chudnovsky and Sivaraman [J.
-Graph Theory 90 (2019) 54-60] proved that every (odd hole, bull)-free graph and every ($P_{5}$, bull)-free graph are perfectly divisible.
-Karthick et al.
-[The Electron.
-J.
-of Combin.
-29 (2022) P3.19.] proved that every (fork, bull)-free graph is perfectly divisible.
+Graph Theory 90 (2019) 54-60] proved (bull, $P_{5}$)-free graphs are perfectly divisible.
 Chen and Xu [Discrete Appl.
 Math.
-372 (2025) 298-307.] proved that every ($P_{7},C_{5}$, bull)-free graph is perfectly divisible.
-Let $H\in${{odd torch}, $\{P_{8},C_{5}\}\}$.
-In this paper, we prove that every ($H$, bull)-free graph is perfectly divisible.
-We also prove that a ($P_{6}$, bull)-free graph is perfectly divisible if and only if it contains no Mycielski-Grötzsch graph as an induced subgraph.
-As corollaries, these graphs are $\binom{\omega+1}{2}-$colorable.
-Notice that every odd torch contains an odd hole, an induced $P_{5}$, and an induced fork.
-Therefore, our results generalize their findings.
-Moreover, we prove that every ($P_{6}$, bull)-free graph $G$ satisfies $\chi(G)\leq\omega(G)^{7}$.
-Key words and phrases: bull-free graphs; perfectly divisibility; chromatic number.
-AMS 2000 Subject Classifications: 05C15, 05C75
+372 (2025) 298-307] proved that (bull, $P_{7},C_{5}$)-free graphs are perfectly divisible.
+In this paper, we prove that (bull, $P_{8},C_{5}$)-free graphs are perfectly divisible.
+As known $F$ is non-perfectly divisible, and so the class of (bull, $P_{6}$)-free graphs is non-perfectly divisible.
+Let $G$ be a (bull, $P_{6}$)-free graph.
+We also show that (i) $G$ is perfectly divisible if and only if $G$ is $F$-free; (ii) if $G$ is odd net-free, then $G$ can be partitioned into two perfect graphs, and so $\chi(G)\leq 2\omega(G)$; (iii) if $G$ is odd net+-free, then $\chi(G)\leq\omega(G)^{2}$; (iv) $\chi(G)\leq\omega(G)^{7}$.
+These results generalize the findings of Chudnovsky and Sivaraman, and Chen and Xu.
+Key words and phrases: bull-free graphs; perfect divisibility; chromatic number.
+AMS 2020 Subject Classifications: 05C15, 05C75
 1 introduction
@@
 Analogously, for a family $\mathcal{H}$ of graphs, we say that $G$ is $\mathcal{H}$-free if $G$ induces no member of $\mathcal{H}$.
-For a given positive integer $k$, we use the notation $[k]$ to denote the set $\{1,\ldots,k\}$.
+Let $H_{1}$ and $H_{2}$ be two vertex disjoint graphs.
+The join $H_{1}+H_{2}$ is the graph with $V(H_{1}+H_{2})=V(H_{1})\cup V(H_{2})$ and $E(H_{1}+H_{2})=E(H_{1})\cup E(H_{2})\cup\{xy\;|\;x\in V(H_{1}),y\in V(H_{2})$}\}.
+For a given positive integer $k$, we use $[k]$ to denote the set $\{1,\ldots,k\}$.
 A k-coloring of a graph $G=(V,E)$ is a mapping $f$: $V\rightarrow[k]$ such that $f(u)\neq f(v)$ whenever $uv\in E$.
@@
 The clique number of $G$, denoted by $\omega(G)$, is the maximum size of a clique in $G$.
-[…]
+The concept of $\chi$-boundedness was introduced by Gyárfás [15] in 1975.
+A class of graphs is said to be hereditary if it is closed under isomorphism and induced subgraphs.
+For a hereditary class $\mathcal{F}$, if there exists a function $f$ such that $\chi(H)\leq f(\omega(H))$ for every $H\in\mathcal{F}$, then $\mathcal{F}$ is called $\chi$-bounded and $f$ a binding function.
 An induced cycle of length $k\geq 4$ is called a hole, and $k$ is the length of the hole.
+A k-hole is a hole of length $k$.
 A hole is odd if $k$ is odd, and even otherwise.
@@
 A graph $G$ is said to be perfect if $\chi(H)=\omega(H)$ for every induced subgraph $H$ of $G$.
-The famous Strong Perfect Graph Theorem [7] was established by Chudnovsky et al.
+The Strong Perfect Graph Theorem [8] was established by Chudnovsky et al.
 in 2006:
@@
 [8] A graph $G$ is perfect if and only if $G$ is ((odd hole, odd antihole))-free.
-Let $F$ denote the Mycielski-Grötzsch graph $G_{4}$.
-A bull is a graph consisting of a triangle with two disjoint pendant edges, an odd torch is a graph obtained from an odd hole by adding an edge $xy$ such that $x$ is non-adjacent to any vertex on the odd hole and the set of neighbors of $y$ on the odd hole is a stable set, a fork is a graph obtained from $K_{1,3}$ by subdividing an edge once, and a E is a graph obtained from an induced $P_{5}=v_{1}v_{2}v_{3}v_{4}v_{5}$ by adding an new vertex exactly adjacent to $v_{3}$ in the $P_{5}$.
-(See Figure 1)
-Figure 1: Illustration of some graphs.
-A graph is perfectly divisible if for each induced subgraph $H$ of $G$, $V(H)$ can be partitioned into $A$ and $B$ such that $H[A]$ is perfect and $\omega(H[B])<\omega(H)$.
-This concept was proposed by Hoáng in [12].
-Chudnovsky and Sivaraman [9] proved that every (odd hole, bull)-free graph and every ($P_{5}$, bull)-free graph are perfectly divisible.
-Karthick et al.
-[15] proved that every (fork, bull)-free graph is perfectly divisible, and this class of graphs was generalized by Wu and Xu [19].
-Chen and Xu [5] proved that every ($P_{7},C_{5}$, bull)-free graph is perfectly divisible.
-In this paper, we prove the following theorems.
-Theorem 1.2
-Let $G$ be a (oddtorch,bull)(odd~torch,bull)-free graph. Then $G$ is perfectly divisible.
-Theorem 1.3
-Let $G$ be a $(P_{8},C_{5},bull)$-free graph. Then $G$ is perfectly divisible.
-It is easy to verify that every odd torch contains an odd hole, an induced $P_{5}$, an induced fork, and the graph $E$.
-Therefore, all the class of (odd hole, bull)-free graphs, ($P_{5}$, bull)-free graphs, (fork, bull)-free, and ($E$, bull)-free graphs are subclasses of the class of (odd torch, bull)-free graphs.
-Theorem 1.2 generalizes the results of Chudnovsky and Sivaraman [9] and Karthick et al.
-[15]; and Theorem 1.3 generalizes the result of Chen and Xu [5].
-Recall that we use $F$ to denote the Mycielski-Grötzsch graph.
-It is certain that $F$ is ($P_{6}$, bull)-free and not perfectly divisible (because $\chi(F)=4$ and $\omega(F)=2$.).
-In fact, the graph $G$ obtained from $F$ by adding a $K_{n}$ in which each vertex is adjacent to all the vertices in $F$ is ($P_{6}$, bull)-free with clique number $n+2$ and not perfectly divisible.
-Since every ($P_{5}$, bull)-free graph is perfectly divisible, a natural problem is that under what conditions is ($P_{6}$, bull)-free graph perfectly divisible ?
-In this paper, we prove the following theorem.
-Theorem 1.4
-Let $G$ be a $(P_{6},\text{bull})$-free graph. Then $G$ is perfectly divisible if and only if $G$ is $F$-free.
-[…]
-By a simple induction on $\omega(G)$, we have that $\chi(G)\leq\binom{\omega(G)+1}{2}$ for each perfectly divisible graph $G$.
-According to Theorems 1.2, 1.3 and 1.4, we can directly derive the following corollary.
-Corollary 1.1
-Let $G$ be an $(H,\text{bull})$-free graph, where $H\in\{\{odd~torch\},\{P_{8},C_{5}\},\{P_{6},F\}\}$ Then $\chi(G)\leq\binom{\omega(G)+1}{2}$.
-Corollary 1.2
-Let $G$ be an ((odd torch, $C_{3}$))-free graph. Then $\chi(G)\leq 3$.
-We also consider the binding function of ($P_{6}$, bull)-free graphs, and prove the following theorem.
-Theorem 1.5
-Let $G$ be a $(P_{6},bull)$-free graph. Then $\chi(G)\leq\omega(G)^{7}$.
+Let $F$ denote the Grötzsch graph.
+A bull is a graph consisting of a triangle with two disjoint pendant edges.
+Let $H$ be an odd hole or an odd antihole.
+An odd net is the graph obtained from $H$ by adding a new vertex adjacent to all vertices of $H$.
+An odd net+ is the graph obtained from an odd net by adding an additional new vertex that is non-adjacent to any vertex of $H$.
+(see Figure 1).
+Figure 1: Illustration of some forbidden graphs.
+A graph $G$ is perfectly divisible if for each induced subgraph $H$ of $G$, $V(H)$ can be partitioned into $A$ and $B$ such that $H[A]$ is perfect and $\omega(H[B])<\omega(H)$.
+This concept was proposed by Hoàng in [13].
+By a simple induction on $\omega(G)$, we have that every perfectly divisible graph $G$ satisfies $\chi(G)\leq\binom{\omega(G)+1}{2}$ [10].
 Let $G$ be a graph and let $w$ be a weight function on $V(G)$.
 We use $\omega_{w}(G)$ to denote the maximum weight of a clique in $G$.
-A graph $G$ is perfectly weight divisible [9, 12] if for every positive integer weight function $w$ on $V(G)$, and for every induced subgraph $H$ of $G$, there is a partition of $V(H)$ into $P$, $W$ such that $H[P]$ is perfect and $\omega_{w}(H[W])<\omega_{w}(H)$.
+A graph $G$ is perfectly weight divisible [10, 13] if for every positive integer weight function $w$ on $V(G)$, and for every induced subgraph $H$ of $G$, there is a partition of $V(H)$ into $P$ and $W$ such that $H[P]$ is perfect and $\omega_{w}(H[W])<\omega_{w}(H)$.
+If $G$ is not perfectly weight divisible but each of its proper induced subgraph is, then we say $G$ is minimally non-perfectly weight divisible (MNWD for short).
 It is certain that if a graph $G$ is perfectly weight divisible, then $G$ is perfectly divisible.
-If $G$ is not perfectly weight divisible but each of its proper induced subgraph is, then we say $G$ is minimally non-perfectly weight divisible (MNWD for short).
-As mentioned above, the structure and coloring problem of bull-free graphs has garnered significant attention, with new methods and research content [4, 6, 11, 13, 18] being continuously proposed.
-This paper utilizes the concept of perfectly weight divisibility to further analyze the structure of bull-free graphs and bull-free MNWD graphs, thereby deriving our main results.
-We will prove Theorem 1.2 in Section 3, prove Theorems 1.4 and 1.3 in Section 4, and prove Theorem 1.5 in Section 5.
+Chudnovsky and Sivaraman [10] proved that every (bull, odd hole)-free graph and every (bull, $P_{5}$)-free graph are perfectly divisible.
+Karthick et al.
+[16] proved that every (bull, fork)-free graph is perfectly divisible, where fork is the graph obtained from a $K_{1,3}$ by subdividing an edge once.
+Chen and Xu [5] proved that every (bull, $P_{7},C_{5}$)-free graph is perfectly divisible.
+It is certain that $F$ is (bull, $P_{6}$)-free and not perfectly divisible (because $\chi(F)=4$ and $\omega(F)=2$).
+Actually, the graph $F+K_{n}$ is ($P_{6}$, bull)-free with clique number $n+2$ and not perfectly divisible as $F$ is not perfectly divisible.
+Further properties of perfectly divisible graphs and bull-free graphs can be found in [6, 12, 14].
+Since every (bull, $P_{5}$)-free graph is perfectly divisible, a natural question is to determine the conditions under which a (bull, $P_{6}$)-free graph is perfectly divisible.
+In this paper, we generalize the results of Chudnovsky and Sivaraman [10], and Chen and Xu [5] by the following two theorems.
+Theorem 1.2
+Let $G$ be a $(bull,P_{8},C_{5})$-free graph. Then $G$ is perfectly weight divisible, and hence $G$ is perfectly divisible.
+Theorem 1.3
+Let $G$ be a $(bull,P_{6})$-free graph. Then $G$ is perfectly divisible if and only if $G$ is $F$-free.
+Note that the class of $3K_{1}$-free graphs has no linear binding function [19]. In 2022, Brause et al.[3] gave a more general theorem.
+Theorem 1.4
+[…]
+By Theorem 1.4, the classes of $(\text{bull},H)$-free graphs, where $H\in\{\{P_{8},C_{5}\},\{P_{6},F\}\}$, do not have a linear binding function.
+By Theorems 1.2 and 1.3, we can directly derive the following corollaries by a simple induction on $\omega(G)$.
+Corollary 1.1
+Let $G$ be an $(\text{bull},H)$-free graph, where $H\in\{\{P_{8},C_{5}\},\{P_{6},F\}\}$ Then $\chi(G)\leq\binom{\omega(G)+1}{2}$.
+[…]
+Theorem 1.5
+Let $G$ be a ((bull, $P_{6}$, odd net))-free graph. Then $G$ can be partitioned into two perfect graphs, and hence $\chi(G)\leq 2\omega(G)$.
+Theorem 1.6
+[…]
+Theorem 1.7
+Let $G$ be a $(bull,P_{6})$-free graph. Then $\chi(G)\leq\omega(G)^{7}$.
+We will prove Theorem 1.2 in Section 3, prove Theorem 1.3 in Section 4, and prove Theorems 1.5, 1.6 and 1.7 in Section 5.
 2 Notations and Preliminary Results
@@
 If it does not cause any confusion, we usually omit the subscript $G$ and simply write $N(v)$, $d(v)$, $M(v)$, $N(X)$ and $M(X)$.
-For a subset $A$ of $V(G)$ and a vertex $b\in V(G)\setminus A$, we say that $b$ is complete to $A$ if $b$ is adjacent to every vertex of $A$, and that $b$ is anticomplete to $A$ if $b$ is not adjacent to any vertex of $A$.
+For a subset $A$ of $V(G)$ and a vertex $u\in V(G)\setminus A$, we say that $u$ is complete to $A$ if $u$ is adjacent to every vertex of $A$, and that $u$ is anticomplete to $A$ if $u$ is not adjacent to any vertex of $A$.
 For two disjoint subsets $A$ and $B$ of $V(G)$, AA is complete to $B$ if every vertex of $A$ is complete to $B$, and $A$ is anticomplete to $B$ if every vertex of $A$ is anticomplete to $B$.
 For $A,B\subseteq V(G)$, let $N_{A}(B)=N(B)\cap A$ and $M_{A}(B)=A\setminus(N_{A}(B)\cup B)$. For $u,v\in V(G)$, we simply write $u\sim v$ if $uv\in E(G)$, and write $u\not\sim v$ if $uv\not\in E(G)$.
-3 Proof of Theorem 1.2
 Let $S\subseteq V(G)$ with $1<|S|<|V(G)|$.
 We say that $S$ is a homogeneous set of $G$ if every vertex in $V(G)\setminus S$ is either complete to $S$ or anticomplete to $S$.
-We say that a graph $G$ is locally perfect if for every vertex $v\in V(G)$, $G[N(v)]$ is perfect.
-In this section, we will prove Theorem 1.2. The following useful lemmas is important to our proof.
+The following useful lemmas are important to our proof.
 Lemma 2.1
-[9] Every MNWD graph has no homogeneous set.
+((Theorem 3.6 of [10])) Every MNWD graph has no homogeneous set.
 Lemma 2.2
-[8] If $G$ is a bull-free graph, then either $G$ has a homogeneous set or for every $v\in V(G)$, either $G[N(v)]$ is perfect or $G[M(v)]$ is perfect.
+((4.3 of [9])) If $G$ is a bull-free graph, then either $G$ has a homogeneous set or for every $v\in V(G)$, either $G[N(v)]$ is perfect or $G[M(v)]$ is perfect.
 Lemma 2.3
-Let $G$ be a connected locally perfect bull-free graph, and let $v\in V(G)$. Then $G[M(v)]$ contains no odd antihole of number of vertices at least 7.
-Proof.
-Suppose to its contrary that $H$ is an odd antihole of $G[M(v)]$ with $V(H)=\{v_{1},v_{2},...,v_{k}\}$, where $k\geq 7$ and $\overline{H}=v_{1}v_{2}\cdots v_{k}v_{1}$.
-Let $v^{\prime}\in N(v)$.
-We first prove that
-if $y\in N(V(H))$, then $|N(y)\cap V(H)|\geq 2$. (1)
-Suppose not, and we may by symmetry assume that $N(y)\cap V(H)=\{v_{1}\}$. But then $\{v_{1},v_{3},v_{5},y,v_{2}\}$ induces a bull, a contradiction. This proves (1).
-We next prove that
-if $x\in M(V(H))$ and $y\in N(V(H))$ such that $x\sim y$, then $N(y)\cap V(H)$ is a stable set. (2)
-Suppose not, and without loss of generality, we may suppose that $v_{1},v_{n}\in N(y)\cap V(H)$ with $v_{1}v_{n}\in E(G)$, where $3\leq n\leq k-2$.
-We next show that $y$ is complete to $\{v_{1},v_{2},\cdots,v_{n}\}$.
-If not, then let $2\leq n^{\prime}\leq n-1$ be the minimum integer such that $y\not\sim v_{n^{\prime}}$.
-If $n=3$, then $n^{\prime}=2$.
-To avoid an induced bull on $\{v_{1},v_{3},y,x,v_{4}\}$, we have that $y\sim v_{4}$; but then $\{v_{1},v_{4},y,x,v_{2}\}$ induces a bull, a contradiction.
-Hence, $n\geq 4$, and thus $v_{n}\sim v_{2}$ and $v_{n-1}\sim v_{1}$.
-We can deduce that $n^{\prime}\neq 2$ to avoid an induced bull on $\{v_{1},v_{n},y,x,v_{2}\}$; and $n^{\prime}\neq n-1$ to avoid an induced bull on $\{v_{1},v_{n},y,x,v_{n-1}\}$.
-We have that $3\leq n^{\prime}\leq n-2$, and so $v_{n^{\prime}}\sim v_{n}$ and $v_{1}\sim v_{n^{\prime}-1}$ by the minimality of $n^{\prime}$.
-But then $\{v_{n^{\prime}-1},v_{n},y,x,v_{n^{\prime}},\}$ induces a bull, a contradiction.
-Therefore, $y$ is complete to $\{v_{1},v_{2},\cdots,v_{n}\}$.
-Similarly, with the same arguments, we can deduce that $y$ is complete to $\{v_{1},v_{k},v_{k-1},\cdots,v_{n}\}$, and this implies that $y$ is complete to $V(H)$, which contradicts to that $G$ is locally perfect.
-This proves (2).
-Since $G$ is connected and $V(H)\subseteq M(v)$, there exist two adjacent vertices $x\in M(V(H))$ and $y\in N(V(H))$.
-By (1) and (2), we may by symmetry assume that $N(y)\cap V(H)=\{v_{1},v_{2}\}$.
-But then, $\{v_{k-3},v_{k-1},v_{k},v_{1},y\}$ induces a bull, a contradiction.
-This completes the proof of Lemma 3.3.
-$\square$
+((Lemma 6 of [4])) Let $G$ be a bull-free MNWD graph.
+Then $G$ is a connected odd net-free graph, and for every vertex $v\in V(G)$, $G[M(v)]$ is imperfect and contains no odd antihole of number of vertices at least $7$.
 Lemma 2.4
-Let $G$ be a bull-free MNWD graph. Then $G$ is a connected locally perfect graph, and for every vertex $v\in V(G)$, $G[M(v)]$ is imperfect.
-Proof.
-It is certain that $G$ is connected.
-Since $G$ is MNWD, we have that for every vertex $v\in V(G)$, $G[M(v)]$ cannot be perfect as otherwise, $(\{v\}\cup M(v),N(v))$ is a partition of $V(G)$ such that for all positive integer weight function on $V(G)$, the $G[\{v\}\cup M(v)]$ is a perfect graph and $G[N(v)]$ has a smaller maximum weight of a clique than $G$, which contradicts with the minimality of $G$.
-Therefore, for every $v\in V(G)$, $G[M(v)]$ is imperfect.
-Moreover, since $G$ is MNWD, $G$ has no homogeneous set by Lemma 3.1.
-Therefore, for every $v\in V(G)$, $G[N(v)]$ is perfect by Lemma 3.2.
-This proves Lemma 3.4.
-$\square$
-By Lemmas 3.3 and 3.4, we have that
+((Theorem 2 of [11])) Let $G$ be a $(P_{8},C_{5},C_{3})$-free graphs. Then $\chi(G)\leq 3$.
 Lemma 2.5
-Let $G$ be a bull-free MNWD graph, and let $v\in V(G)$. Then $G[M(v)]$ contains no odd antihole of number of vertices at least 7.
-We now prove Theorem 1.2.
-Proof of Theorems 1.2: It suffices to show that each (odd torch, bull)-free graph is perfectly weight divisible.
-Suppose it is not true, and let $G$ be a (odd torch, bull)-free MNWD graph.
-By Lemma 3.4, we have that $G$ is connected, and
-for every vertex $v\in V(G)$, $G[N(v)]$ is perfect and $G[M(v)]$ is imperfect. (3)
-Let $v\in V(G)$.
-By (3), Lemma 3.5 and Theorem 1.1, $G[M(v)]$ contains an odd hole, say $C=v_{1}v_{2}\cdots v_{k}v_{1}$, where $k\geq 5$ is an odd integer.
-Since $G$ is connected, there exist two adjacent vertices $x\in M(V(C))$ and $y\in N(V(C))$.
-Since $G$ is odd torch-free, we have that $N(y)\cap V(C)$ is not a stable set.
-Hence, we may by symmetry assume that $\{v_{1},v_{2}\}\subseteq N(y)\cap V(C)$.
-By (3), we have that $y$ is not complete to $V(C)$ as $G[N(y)]$ is perfect.
-Let $3\leq n\leq k$ be the minimum integer such that $y\not\sim v_{n}$.
-But then $\{v_{n-2},v_{n-1},y,x,v_{n}\}$ induces a bull, a contradiction.
-Therefore, each (odd torch, bull)-free graph is perfectly weight divisible.
-This completes the proof of Theorem 1.2.
-$\square$
-4 Proof of Theorems 1.3 and 1.4
-In this section, we will prove Theorems 1.3 and 1.4. The following useful lemmas are important to our proof.
+((Theorem 6 of [18])) Let $G$ be a $(P_{6},C_{3})$-free graphs. Then $\chi(G)\leq 4$, and $G$ contains $F$ if $\chi(G)=4$.
 Lemma 2.6
@@
 Since $\chi(G^{\prime})\leq 3$, there is a partition $(A,B)$ of $V(G^{\prime})$ such that $G^{\prime}[A]$ is bipartite and $B$ is a stable set of $G^{\prime}$.
-Let $X\subseteq B$ such that each vertex in $X$ is a isolated vertex of $G^{\prime}$.
+Let $X\subseteq B$ such that each vertex in $X$ is an isolated vertex of $G^{\prime}$.
 Then $(A\cup X,B\setminus X)$ is a partition of $G^{\prime}$ such that $G[A\cup X]$ is perfect, and $\omega_{w}(G[B\setminus X])<\omega_{w}(G^{\prime})$.
@@
 $\square$
+3 Proof of Theorem 1.2
+In this section, we prove Theorem 1.2 and begin with the following lemma.
 Lemma 3.1
-Let $G$ be a connected locally perfect $(P_{6},bull)$-free graph with $\omega(G)\geq 3$.
-Let $v\in V(G)$ such that $v$ belongs to a maximum clique of $G$.
-If $G[M(v)]$ contains a 5-hole, then $G$ has a homogeneous set.
-Proof. Let $C=v_{1}v_{2}\cdots v_{5}v_{1}$ be a 5-hole in $G[M(v)]$. From now on, the subscript is modulo 5 in the proof of the Lemma. For $1\leq i\leq 5$, let
-$X_{i}$ $=$ $\{u\in N(V(C))~|~N_{C}(u)=\{v_{i}\}\};$
-$Y_{i}$ $=$ $\{u\in N(V(C))~|~N_{C}(u)=\{v_{i},v_{i+2}\}\};$
-$Z_{i}$ $=$ $\{u\in N(V(C))~|~N_{C}(u)=\{v_{i},v_{i+1},v_{i+2}\}\};$
-$W_{i}$ $=$ $\{u\in N(V(C))~|~N_{C}(u)=\{v_{i},v_{i+1},v_{i+2},v_{i+3}\}\};$
-Let $X=\bigcup_{i=1}^{5}X_{i}$, $Y=\bigcup_{i=1}^{5}Y_{i}$, $Z=\bigcup_{i=1}^{5}Z_{i}$ and $W=\bigcup_{i=1}^{5}W_{i}$. We next prove the following claim.
-Claim 4.1
-$N(V(C))=X\cup Y\cup Z\cup W$.
-Proof.
-It suffices to prove that $N(V(C))\subseteq X\cup Y\cup Z\cup W$.
-Let $u\in N(V(C))$.
-Without loss of generality, suppose $u\sim v_{1}$.
-If $N(u)\cap V(C)=\{v_{1}\}$, then $u\in X_{1}$, and we are done.
-So, suppose $u$ has a neighbor in $\{v_{2},v_{3},v_{4},v_{5}\}$.
-If $u$ is anticomplete to $\{v_{2},v_{5}\}$, then $x$ is not complete to $\{v_{3},v_{4}\}$ as otherwise $\{u,v_{3},v_{4},v_{2},v_{5}\}$ induces a bull.
-Hence, $u$ has exactly one neighbor in $\{v_{3},v_{4}\}$.
-Now, $u\in Y_{1}$ if $u\sim v_{3}$ and $u\in Y_{4}$ if $u\sim v_{4}$.
-Therefore, we may assume that $u$ has a neighbor in $\{v_{2},v_{5}\}$, and by symmetry assume that $u\sim v_{2}$.
-To forbid an induced bull on $\{u,v_{1},v_{2},v_{3},v_{5}\}$, $u$ has a neighbor in $\{v_{3},v_{5}\}$.
-By symmetry, we may assume that $u\sim v_{3}$.
-If $N(u)\cap V(C)=\{v_{1},v_{2},v_{3}\}$, then $u\in Z_{1}$, and we are done.
-So, we may suppose that $u$ has a neighbor in $\{v_{4},v_{5}\}$.
-Since $G$ is locally perfect, we have that $u$ is not complete to $V(C)$.
-Now, $u\in W_{1}$ if $u\sim v_{4}$ and $u\in W_{5}$ if $u\sim v_{5}$.
-This proves Claim 4.1.
-$\square$
-Consequently, we prove that
-for every $u\in N(v)$, $u\in Y$. (4)
-Since $G$ is connected, we claim that $u$ has a neighbor in $V(C)$.
-On the contrary, there exists an induced $P_{3}=x_{1}x_{2}x_{3}$ such that $\{x_{1},x_{2}\}\subseteq M(V(C))$ and $x_{3}\in N(V(C))$.
-By symmetry, suppose $x_{3}\sim v_{1}$.
-By Claim 4.1, without loss of generality, we may assume that $x_{3}\in X_{1}\cup Y_{1}\cup Z_{1}\cup W_{1}$.
-If $x_{3}\in W_{1}$, then $\{x_{3},v_{3},v_{4},x_{2},v_{5}\}$ induces a bull; if $x_{3}\in X_{1}$, then there is an induced $P_{6}=x_{1}x_{2}x_{3}v_{1}v_{2}v_{3}$; otherwise, there is an induced $P_{6}=x_{1}x_{2}x_{3}v_{3}v_{4}v_{5}$.
-All are contradictions.
-This implies that $N(u)\cap V(C)\neq\mbox{{\rm\O }}$.
-By Claim 4.1, without loss of generality, we assume that $u\in X_{1}\cup Y_{1}\cup Z_{1}\cup W_{1}$.
-If $u\in X_{1}$, then there is an induced $P_{6}=vuv_{1}v_{2}v_{3}v_{4}$.
-If $u\in Z_{1}$, then $\{v,u,v_{2},v_{3},v_{4}\}$ induces a bull.
-If $u\in W_{1}$, then $\{u,v_{3},v_{4},v,v_{5}\}$ induces a bull.
-So, $u\in Y_{1}$.
-This proves (4).
-We next prove that
-if $x,y\in N(v)$ such that $x\sim y$, then $N_{C}(x)=N_{C}(y)$. (5)
-Assume for contradiction that $N_{C}(x)\neq N_{C}(y)$.
-Without loss of generality, suppose $N_{C}(x)=\{v_{1},v_{3}\}$ by (4).
-If $N_{C}(y)=\{v_{2},v_{4}\}$, then $\{v,x,y,v_{1},v_{4}\}$ induces a bull.
-If $N_{C}(y)=\{v_{3},v_{5}\}$, then $\{v_{3},x,y,v_{2},v_{5}\}$ induces a bull.
-By symmetry, in all other cases an induced bull also occurs.
-Hence, $N_{C}(x)=N_{C}(y)$.
-This proves (5).
-Since $\omega(G)\geq 3$ and $v$ belongs to a maximum clique of $G$, there exist two adjacent vertices in $N(v)$.
-Let $S$ be a component of $N(v)$ which contains an edge.
-It is certain that $1<|V(S)|<|V(G)|$.
-We next prove that $V(S)$ is a homogeneous set of $G$.
-By (4) and (5), we may by symmetry assume that $N(V(S))\cap V(C)=\{v_{1},v_{3}\}$ as $S$ is connected.
-Obviously, we have that $\{v,v_{1},v_{3}\}$ is complete to $V(S)$, and $(N(v)\setminus V(S))\cup\{v_{2},v_{4},v_{5}\}$ is anticomplete to $V(S)$.
-Suppose that $V(S)$ is not a homogeneous set of $G$.
-There exists a vertex $z$ which has a neighbor and a non-neighbor in $V(S)$.
-Clearly, $z\notin N(v)\cup\{v\}\cup V(C)$.
-Since $S$ is connected, there exist an edge $xy\in G[S]$ such that $z\sim x$ and $z\not\sim y$.
-To avoid an induced bull on $\{x,y,z,v_{1},v_{2}\}$, we have that $z\sim v_{2}$.
-Similarly, we have that $z$ is complete to $\{v_{2},v_{4},v_{5}\}$.
-So $z\in N(V(C))$.
-By Claim 4.1, $z$ has a neighbor in $\{v_{1},v_{3}\}$.
-By symmetry, we may assume that $z\sim v_{1}$.
-But then $\{v_{1},x,z,v,v_{4}\}$ induces a bull, a contradiction.
-Therefore, $V(S)$ is a homogeneous set of $G$.
-This completes the proof of Lemma 4.2.
-$\square$
-The proof of the following lemma is similar to that of Lemma 4.2.
-Lemma 4.3
-Let $G$ be a connected locally perfect $(P_{8},C_{5},bull)$-free graph with $\omega(G)\geq 3$.
-Let $v\in V(G)$ such that $v$ belongs to a maximum clique of $G$.
+Let $G$ be a connected $(bull,P_{8},C_{5},odd~net)$-free graph with $\omega(G)\geq 3$, and let $v\in V(G)$ such that $v$ belongs to a maximum clique of $G$.
 If $G[M(v)]$ contains a $7$-hole, then $G$ has a homogeneous set.
-Proof. Let $C=v_{1}v_{2}\cdots v_{5}v_{6}v_{7}v_{1}$ be a 7-hole in $G[M(v)]$. For the remainder of the proof, all subscripts are taken modulo 7. For $1\leq i\leq 7$, let
+Proof. Let $C=v_{1}v_{2}\cdots v_{5}v_{6}v_{7}v_{1}$ be a 7-hole in $G[M(v)]$. For the remainder of the proof, all subscripts are taken modulo 7. Let $i\in[7]$, define
 $X_{i}$ $=$ $\{u\in N(V(C))~|~N_{C}(u)=\{v_{i}\}\};$
@@
 Let $u\in N(V(C))$.
-Without loss of generality, suppose $u\sim v_{1}$.
-If $N(u)\cap V(C)=\{v_{1}\}$, then $u\in X_{1}$, and we are done.
-So, suppose $u$ has a neighbor in $\{v_{2},v_{3},\cdots,v_{7}\}$.
+Without loss of generality, suppose that $u\sim v_{1}$.
+If $N(u)\cap V(C)=\{v_{1}\}$, then $u\in X_{1}$ and we are done.
+So we may assume that $u$ has a neighbor in $\{v_{2},v_{3},\cdots,v_{7}\}$.
 Suppose that $u$ is anticomplete to $\{v_{2},v_{7}\}$.
@@
 Assume that $u$ is anticomplete to $\{v_{3},v_{6}\}$.
-If $u\sim v_{4}$, then there is a 5-hole $uv_{1}v_{2}v_{3}v_{4}u$, and we are done by Lemma 4.2.
-So, $u\not\sim v_{4}$; similarly, $u\not\sim v_{5}$, a contradiction.
+If $u\sim v_{4}$, then there is a 5-hole $uv_{1}v_{2}v_{3}v_{4}u$, a contradiction.
+So, $u\not\sim v_{4}$; similarly, $u\not\sim v_{5}$.
 Therefore, $u$ is not anticomplete to $\{v_{3},v_{6}\}$ and we may by symmetry assume that $u\sim v_{3}$.
-If $u\sim v_{6}$, then $u$ has a neighbor in $\{v_{4},v_{5}\}$ to avoid a 5-hole $uv_{3}v_{4}v_{5}v_{6}u$; and we may by symmetry assume that $u\sim v_{4}$, but now $\{v_{3},v_{4},u,v_{2},v_{6}\}$ induces a bull.
+If $u\sim v_{6}$, then $u$ has a neighbor in $\{v_{4},v_{5}\}$ to avoid a 5-hole $uv_{3}v_{4}v_{5}v_{6}u$; and we may by symmetry assume that $u\sim v_{4}$, but now $\{v_{4},u,v_{3},v_{2},v_{6}\}$ induces a bull.
 So, $u\not\sim v_{6}$.
 To avoid a 5-hole $uv_{5}v_{6}v_{7}v_{1}u$, we have that $u\not\sim v_{5}$.
-Moreover, to avoid an induced bull on $\{v_{3},v_{4},u,v_{2},v_{5}\}$, we have that $u\not\sim v_{4}$.
-Now, $N(u)\cap V(C)=\{v_{1},v_{3}\}$, then $u\in Y_{1}$ and we are done.
-So, $u$ is not anticomplete to $\{v_{2},v_{7}\}$.
+Moreover, to avoid an induced bull on $\{u,v_{3},v_{4},v_{2},v_{5}\}$, we have that $u\not\sim v_{4}$.
+Now, $N(u)\cap V(C)=\{v_{1},v_{3}\}$, and then $u\in Y_{1}$.
+So, we may assume that $u$ is not anticomplete to $\{v_{2},v_{7}\}$.
 Without loss of generality, $u\sim v_{2}$.
-To avoid an induced bull on $\{v_{1},v_{2},u,v_{3},v_{7}\}$, we may by symmetry assume that $u\sim v_{3}$.
-If $u\sim v_{4}$, then $u\sim v_{5}$ to forbid an induced bull on $\{v_{3},v_{4},u,v_{1},v_{5}\}$; similarly, $u\sim v_{7}$.
-Since $G$ is locally perfect, $u$ is not complete to $V(C)$, and so $u\not\sim v_{6}$; but then $\{v_{4},v_{5},u,v_{2},v_{6}\}$ induces a bull, a contradiction.
+To avoid an induced bull on $\{u,v_{1},v_{2},v_{3},v_{7}\}$, we may assume that $u\sim v_{3}$.
+If $u\sim v_{4}$, then $u\sim v_{5}$ to forbid an induced bull on $\{v_{3},u,v_{4},v_{1},v_{5}\}$; similarly, $u\sim v_{7}$.
+Since $G$ is odd net-free, $u$ is not complete to $V(C)$, and so $u\not\sim v_{6}$; but then $\{v_{4},v_{5},u,v_{2},v_{6}\}$ induces a bull, a contradiction.
 So $u\not\sim v_{4}$, and similarly, $u\not\sim v_{7}$.
-If $u$ is complete to $\{v_{5},v_{6}\}$, then $\{v_{5},v_{6},u,v_{4},v_{7}\}$ induces a bull.
+If $u$ is complete to $\{v_{5},v_{6}\}$, then $\{u,v_{5},v_{6},v_{4},v_{7}\}$ induces a bull.
 So $u$ has at most one neighbor in $\{v_{5},v_{6}\}$.
-If $u\sim v_{5}$, then there is a 5-hole $uv_{5}v_{6}v_{7}v_{1}u$.
-Hence, $u$ is anticomplete to $\{v_{5},v_{6}\}$ by symmetry.
-And thus $N(u)\cap V(C)=\{v_{1},v_{2},v_{3}\}$; this implies $u\in Z_{1}$.
+If $u\sim v_{5}$, then there exists a 5-hole $uv_{5}v_{6}v_{7}v_{1}u$.
+So, $u\not\sim v_{5}$.
+Similarly, $u\not\sim v_{6}$.
+Then $N(u)\cap V(C)=\{v_{1},v_{2},v_{3}\}$; this implies $u\in Z_{1}$.
 This proves Claim 3.1.
@@
 Claim 3.2
-For every $u\in N(v)$, we have that $u\in Y$. And for $x,y\in N(v)$ with $x\sim y$, $N_{C}(x)=N_{C}(y)$.
+If there exists $u\in N(v)$ such that $u\notin N(V(C))$, then $G$ has a homogeneous set.
+Proof.
+Let $u\in N(v)$ such that $u\notin N(V(C))$.
+Then $u\in M(V(C))$.
+Since $G$ is connected, it follows that there exists an induced $P=x_{1}x_{2}\cdots x_{h}$ connecting $uv$ and $C$, where $h\geq 3$, such that $x_{1}\in\{u,v\}$ and $x_{h}\in V(C)$.
+Without loss of generality, $x_{1}=u$ and $x_{h}=v_{1}$.
+Since $x_{h-1}\in N(V(C))$, we may by symmetry assume that $x_{h-1}\in X_{1}\cup Y_{1}\cup Z_{1}$ by Claim 3.1.
+If $x_{h-1}\in X_{1}$, then there exists an induced $P_{8}=x_{h-2}x_{h-1}v_{1}v_{2}v_{3}v_{4}v_{5}v_{6}$.
+If $x_{h-1}\in Z_{1}$, then $\{v_{2},v_{1},x_{h-1},v_{7},x_{h-2}\}$ induces a bull.
+Therefore $x_{h-1}\in Y_{1}$.
+Given that $h=3$ as otherwise there exists an induced $P_{8}=v_{7}v_{6}v_{5}v_{4}v_{3}x_{h-1}x_{h-2}x_{h-3}$.
+So $x_{2}\sim v_{1}$ and $x_{2}\sim v_{3}$ as $x_{h-1}\in Y_{1}$.
+Consequently, to avoid an induced $P_{8}=v_{7}v_{6}v_{5}v_{4}v_{3}x_{2}x_{1}v$, we have that $v\sim x_{2}$.
+Let $S$ be the component of $G[N(x_{2})\cap M(V(C))]$ which contains the edge $uv$.
+We claim that $V(S)$ is a homogeneous set of $G$.
+Suppose that there exists a vertex $t\in V(G)\setminus V(S)$ such that $t$ is neither complete nor anticomplete to $V(S)$.
+So there exist two vertices $t_{1},t_{2}\in V(S)$ such that $tt_{1}t_{2}$ is an induced $P_{3}$.
+Clearly $t\notin V(C)$ as $t_{1}\in M(V(C))$.
+We next show that
+$t\not\sim x_{2}.$ (1)
+Suppose to the contrary that $t\sim x_{2}$.
+By the choice of $S$, we have that $t\in N(V(C))$ as $t\notin V(S)$.
+Then $t$ has at least two neighbors in $V(C)$ because $t$ has a neighbor in $M(V(C))$ and $G$ is $P_{8}$-free.
+Hence $t\in Y\cup Z$ by Claim 3.1.
+Suppose that $t\not\sim v_{1}$.
+Moreover, $t$ is anticomplete to $V(C)\setminus\{v_{2},v_{3},v_{7}\}$ as otherwise $\{t,t_{1},x_{2},v_{1},t^{\prime}\}$ induces a bull, where $t^{\prime}\in V(C)\setminus\{v_{1},v_{2},v_{3},v_{7}\}$ is a neighbor of $t$, a contradiction.
+If $t\sim v_{3}$, then $\{t,x_{2},v_{3},t_{2},v_{2}\}$ induces a bull whenever $t\not\sim v_{2}$, and $\{t,v_{2},v_{3},v_{1},v_{4}\}$ induces a bull whenever $t\sim v_{2}$.
+Both are contradictions.
+So $t\not\sim v_{3}$ and $t$ is anticomplete to $V(C)\setminus\{v_{2},v_{7}\}$.
+By Claim 3.1, $t\in Y_{7}$ as $t$ has at least two neighbors in $V(C)$.
+But then there exists an induced $P_{8}=v_{6}v_{5}v_{4}v_{3}v_{2}tt_{1}t_{2}$.
+So $t\sim v_{1}$, and by symmetry $t\sim v_{3}$.
+Now $t\in Y_{1}\cup Z_{1}$ by Claim 3.1. But then $\{t,x_{2},v_{3},v_{4},t_{2}\}$ induces a bull. This proves (1).
+By (1), we have that $t\sim v_{3}$ to avoid a bull on $\{t_{2},t_{1},x_{2},t,v_{3}\}$.
+Similarly, $t\sim v_{1}$.
+By Claim 3.1, $t\in Y_{1}\cup Z_{1}$.
+But then there exists an induced $P_{8}=v_{7}v_{6}v_{5}v_{4}v_{3}tt_{1}t_{2}$, a contradiction.
+Therefore $V(S)$ is a homogeneous set of $G$.
+This proves Claim 3.2.
+$\square$
+By Claim 3.2, we may assume that
+for every $u\in N(v)$, $u\in N(V(C))$. (2)
+Claim 3.3
+For every $u\in N(v)$, $u\in Y$. And for any $x,y\in N(v)$ with $x\sim y$, $N_{C}(x)=N_{C}(y)$.
 Proof.
 Let $u\in N(v)$.
-If $N_{C}(u)=\mbox{{\rm\O }}$, since $G$ is connected, there exists an induced $P_{3}=x_{1}x_{2}x_{3}$ such that $\{x_{1},x_{2}\}$ is anticomplete to $V(C)$ and $x_{3}$ has a neighbor in $V(C)$.
-We may by symmetry assume that $x_{3}\in X_{3}\cup Y_{1}\cup Z_{1}$ by Claim 4.2; but then there exists an induced $P_{8}=x_{!}x_{2}x_{3}v_{3}v_{4}v_{5}v_{6}v_{7}$.
+Recall that $G$ is connected.
+If $N_{C}(u)=\mbox{{\rm\O }}$, then there exists an induced $P_{3}=x_{1}x_{2}x_{3}$ such that $\{x_{1},x_{2}\}$ is anticomplete to $V(C)$ and $x_{3}$ has a neighbor in $V(C)$.
+We may by symmetry assume that $x_{3}\in X_{1}\cup Y_{1}\cup Z_{1}$ by Claim 3.1; but then there exists an induced $P_{8}=x_{1}x_{2}x_{3}v_{1}v_{7}v_{6}v_{5}v_{4}$.
 So $u\in N(V(C))$.
-By symmetry and Claim 4.2, we may assume that $u\in X_{3}\cup Y_{1}\cup Z_{1}$.
+Without loss of generality, we may assume that $u\in X_{1}\cup Y_{1}\cup Z_{1}$.
 If $u\in X_{1}$, then there is an induced $P_{8}=vuv_{1}v_{2}v_{3}v_{4}v_{5}v_{6}$.
@@
 Clearly, $z\notin N(v)\cup\{v\}\cup V(C)$.
-Since $S$ is connected, there exist an edge $xy\in G[S]$ such that $z\sim x$ and $z\not\sim y$.
+Since $S$ is connected, there exists an edge $xy\in G[S]$ such that $z\sim x$ and $z\not\sim y$.
 To avoid an induced bull on $\{x,y,v_{1},v_{2},z\}$, we have that $z\sim v_{2}$.
@@
 $\square$
-We need the following lemmas.
-Lemma 4.4
-[17] Let $G$ be a $(P_{6},C_{3})$-free graphs. Then $\chi(G)\leq 4$, and $G$ contains Mycielski-Grötzsch graph if $\chi(G)=4$.
-Lemma 4.5
-[10] Let $G$ be a $(P_{8},C_{5},C_{3})$-free graphs. Then $\chi(G)\leq 3$.
-Now, we can prove Theorems 1.3 and 1.4.
-Proof of Theorems 1.3 and 1.4: Since the graph $F$ is non-perfectly divisible, it suffices to prove that every ($P_{8},C_{5}$, bull)-free graph and every ($P_{6}$, bull, $F$)-free graph are perfectly divisible.
-We will actually prove the following stronger conclusion: every ($P_{8},C_{5}$, bull)-free graph and every ($P_{6}$, bull, $F$)-free graph are perfectly weight divisible.
-Suppose not, and let $G$ be a ($H$, bull)-free MNWD graph, where $H\in\{P_{6},\{P_{8},C_{5}\}\}$.
-By Lemma 3.4, we have that $G$ is a connected locally perfect graph, and for every vertex $v\in V(G)$, $G[M(v)]$ is imperfect.
-By Lemmas 4.14.4 and 4.5, we have that $\omega(G)\geq 3$ because $G$ is non-perfectly weight divisible.
+We now proceed to prove Theorem 1.2.
+Proof of Theorem 1.2: On the contrary, let $G$ be a (bull, $P_{8},C_{5}$)-free MNWD graph.
+By Lemma 2.3, we have that $G$ is a connected odd net-free graph, and for every vertex $v\in V(G)$, $G[M(v)]$ is imperfect and contains no odd antihole of number of vertices at least 7.
+By Lemmas 2.4 and 2.6, we have that $\omega(G)\geq 3$ because $G$ is non-perfectly weight divisible.
 Let $v\in V(G)$ such that $v$ belongs to a maximum clique of $G$.
-By Lemma 3.5, $G[M(v)]$ contains no odd antihole of number of vertices at least 7.
-By Theorem 1.1, we have that $G[M(v)]$ contains an odd hole.
-Therefore, by Lemmas 4.2 and 4.3, $G$ has a homogeneous set, which contradicts with Lemma 3.1.
-This completes the proof of Theorems 1.3 and 1.4.
+By Theorem 1.1, we have that $G[M(v)]$ contains a 7-hole as $G$ is ($P_{8},C_{5}$)-free.
+Therefore, by Lemma 3.1, $G$ has a homogeneous set, which contradicts Lemma 2.1.
+This completes the proof of Theorem 1.2.
 $\square$
 4 Proof of Theorem 1.3
-In this section, we will prove Theorem 1.5.
-A class of graphs is hereditary if it is closed under taking induced subgraphs.
+In this section, we will prove Theorem 1.3. The following useful lemma is important to our proof, and its proof is similar to the proof of Lemma 3.1.
+Lemma 4.1
+Let $G$ be a connected $(bull,P_{6},odd~net)$-free graph with $\omega(G)\geq 3$, and let $v\in V(G)$ such that $v$ belongs to a maximum clique of $G$.
+If $G[M(v)]$ contains a 5-hole, then $G$ has a homogeneous set.
+Proof. Let $C=v_{1}v_{2}\cdots v_{5}v_{1}$ be a 5-hole in $G[M(v)]$. From now on, the subscript is modulo 5 in the proof of the lemma. For $i\in[5]$, let
+$X_{i}$ $=$ $\{u\in N(V(C))~|~N_{C}(u)=\{v_{i}\}\};$
+$Y_{i}$ $=$ $\{u\in N(V(C))~|~N_{C}(u)=\{v_{i},v_{i+2}\}\};$
+$Z_{i}$ $=$ $\{u\in N(V(C))~|~N_{C}(u)=\{v_{i},v_{i+1},v_{i+2}\}\};$
+$W_{i}$ $=$ $\{u\in N(V(C))~|~N_{C}(u)=\{v_{i},v_{i+1},v_{i+2},v_{i+3}\}\};$
+Let $X=\bigcup_{i=1}^{5}X_{i}$, $Y=\bigcup_{i=1}^{5}Y_{i}$, $Z=\bigcup_{i=1}^{5}Z_{i}$ and $W=\bigcup_{i=1}^{5}W_{i}$. We next prove the following claim.
+Claim 4.1
+$N(V(C))=X\cup Y\cup Z\cup W$.
+Proof.
+It suffices to prove that $N(V(C))\subseteq X\cup Y\cup Z\cup W$.
+Let $u\in N(V(C))$.
+Without loss of generality, suppose that $u\sim v_{1}$.
+If $N(u)\cap V(C)=\{v_{1}\}$, then $u\in X_{1}$ and we are done.
+So, suppose that $u$ has a neighbor in $\{v_{2},v_{3},v_{4},v_{5}\}$.
+If $u$ is anticomplete to $\{v_{2},v_{5}\}$, then $x$ is not complete to $\{v_{3},v_{4}\}$ as otherwise $\{u,v_{3},v_{4},v_{2},v_{5}\}$ induces a bull.
+Hence, $u$ has exactly one neighbor in $\{v_{3},v_{4}\}$.
+Now, $u\in Y_{1}$ if $u\sim v_{3}$ and $u\in Y_{4}$ if $u\sim v_{4}$.
+Therefore, we may assume that $u$ has a neighbor in $\{v_{2},v_{5}\}$, and by symmetry assume that $u\sim v_{2}$.
+To forbid an induced bull on $\{u,v_{1},v_{2},v_{3},v_{5}\}$, $u$ has a neighbor in $\{v_{3},v_{5}\}$.
+By symmetry, we may assume that $u\sim v_{3}$.
+If $N(u)\cap V(C)=\{v_{1},v_{2},v_{3}\}$, then $u\in Z_{1}$ and we are done.
+So, we may suppose that $u$ has a neighbor in $\{v_{4},v_{5}\}$.
+Since $G$ is odd net-free, we have that $u$ is not complete to $V(C)$.
+Now, $u\in W_{1}$ if $u\sim v_{4}$ and $u\in W_{5}$ if $u\sim v_{5}$.
+This proves Claim 4.1.
+$\square$
+We next prove that
+Claim 4.2
+If there exists $u\in N(v)$ such that $u\notin N(V(C))$, then $G$ has a homogeneous set.
+Proof.
+Let $u\in N(v)$ such that $u\notin N(V(C))$.
+Then $u\in M(V(C))$.
+Since $G$ is connected, it follows that there exists an induced $P=x_{1}x_{2}\cdots x_{h}$ connecting $uv$ and $C$, where $h\geq 3$, such that $x_{1}\in\{u,v\}$ and $x_{h}\in V(C)$.
+Without loss of generality, $x_{1}=u$ and $x_{h}=v_{1}$.
+Since $x_{h-1}\in N(V(C))$, we may by symmetry assume that $x_{h-1}\in X_{1}\cup Y_{1}\cup Z_{1}\cup W_{1}$ by Claim 4.1.
+If $x_{h-1}\in X_{1}$, then there exists an induced $P_{6}=x_{h-2}x_{h-1}v_{1}v_{2}v_{3}v_{4}$.
+If $x_{h-1}\in Z_{1}\cup W_{1}$, then $\{v_{2},v_{1},x_{h-1},v_{5},x_{h-2}\}$ induces a bull.
+Therefore $x_{h-1}\in Y_{1}$.
+Given that $h=3$ as otherwise there exists an induced $P_{6}=v_{5}v_{4}v_{3}x_{h-1}x_{h-2}x_{h-3}$.
+So $x_{2}\sim v_{1}$ and $x_{2}\sim v_{3}$ as $x_{h-1}\in Y_{1}$.
+Consequently, to avoid an induced $P_{6}=v_{5}v_{4}v_{3}x_{2}x_{1}v$, we have that $v\sim x_{2}$.
+Let $S$ be the component of $G[N(x_{2})\cap M(V(C))]$ which contains the edge $uv$.
+We claim that $V(S)$ is a homogeneous set of $G$.
+Suppose to the contrary that there exists a vertex $t\in V(G)\setminus V(S)$ such that $t$ is neither complete nor anticomplete to $V(S)$.
+So there exist two vertices $t_{1},t_{2}\in V(S)$ such that $tt_{1}t_{2}$ is an induced $P_{3}$.
+Clearly $t\notin V(C)$ as $t_{1}\in M(V(C))$.
+We next show that
+$t\not\sim x_{2}.$ (3)
+Suppose to the contrary that $t\sim x_{2}$.
+By the choice of $S$, we have that $t\in N(V(C))$ as $t\notin V(S)$.
+It is certain that $t$ has at least two neighbors in $V(C)$ because $t$ has a neighbor in $M(V(C))$ and $G$ is $P_{6}$-free.
+Suppose that $t\not\sim v_{1}$.
+Then $t\not\sim v_{4}$ as otherwise $\{t,t_{1},x_{2},v_{1},v_{4}\}$ induces a bull, a contradiction.
+Consequently, $t\not\sim v_{3}$ as otherwise $\{t,x_{2},v_{3},t_{2},v_{2}\}$ induces a bull whenever $t\not\sim v_{2}$, and $\{t,v_{2},v_{3},v_{1},v_{4}\}$ induces a bull whenever $t\sim v_{2}$.
+Both are contradictions.
+So $t$ is anticomplete to $\{v_{1},v_{3},v_{4}\}$.
+By Claim 4.1, $t\in Y_{5}$.
+But now there is an induced $P_{6}=v_{4}v_{3}v_{2}tt_{1}t_{2}$.
+Therefore $t\sim v_{1}$, and by symmetry $t\sim v_{3}$.
+Now, $t\sim v_{2}$ as otherwise $\{t,v_{3},x_{2},v_{2},t_{2}\}$ induces a bull.
+By Claim 4.1, we may by symmetry assume that $t\in Z_{1}\cup W_{1}$.
+But then $\{v_{2},v_{1},t,v_{5},t_{1}\}$ induces a bull.
+This proves (3).
+By (3), we have that $t\sim v_{3}$ to avoid a bull on $\{t_{1},t_{2},x_{2},t,v_{3}\}$.
+Similarly, $t\sim v_{1}$.
+By Claim 4.1, $t\in Y_{1}\cup W_{3}$.
+But then $v_{5}v_{4}v_{3}tt_{1}t_{2}$ is an induced $P_{6}$ if $t\in Y_{1}$, and $\{v_{5},v_{1},t,v_{2},t_{1}\}$ induces a bull if $t\in W_{3}$.
+Both are contradictions.
+Therefore $V(S)$ is a homogeneous set of $G$.
+This proves Claim 4.2.
+$\square$
+By Claim 4.2, we may assume that for every $u\in N(v)$,
+$u\in N(V(C)).$
+Consequently, we prove that
+for every $u\in N(v)$, $u\in Y$. (4)
+By Claim 4.1, without loss of generality, we assume that $u\in X_{1}\cup Y_{1}\cup Z_{1}\cup W_{1}$.
+If $u\in X_{1}$, then there is an induced $P_{6}=vuv_{1}v_{2}v_{3}v_{4}$.
+If $u\in Z_{1}$, then $\{v_{2},u,v_{3},v_{4},v\}$ induces a bull.
+If $u\in W_{1}$, then $\{v_{3},v_{4},u,v,v_{5}\}$ induces a bull.
+So, $u\in Y_{1}$.
+This proves (4).
+We next prove that
+if $x,y\in N(v)$ such that $x\sim y$, then $N_{C}(x)=N_{C}(y)$. (5)
+Assume for contradiction that $N_{C}(x)\neq N_{C}(y)$.
+Without loss of generality, suppose that $N_{C}(x)=\{v_{1},v_{3}\}$ by (4).
+If $N_{C}(y)=\{v_{2},v_{4}\}$, then $\{v,x,y,v_{1},v_{4}\}$ induces a bull.
+If $N_{C}(y)=\{v_{3},v_{5}\}$, then $\{v_{3},x,y,v_{2},v_{5}\}$ induces a bull.
+By symmetry, in all other cases, an induced bull also occurs.
+Hence, $N_{C}(x)=N_{C}(y)$.
+This proves (5).
+Since $\omega(G)\geq 3$ and $v$ belongs to a maximum clique of $G$, there exist two adjacent vertices in $N(v)$.
+Let $S$ be a component of $N(v)$ which contains an edge.
+It is certain that $1<|V(S)|<|V(G)|$.
+We next prove that $V(S)$ is a homogeneous set of $G$.
+By (4) and (5), we may by symmetry assume that $N(V(S))\cap V(C)=\{v_{1},v_{3}\}$ as $S$ is connected.
+It is certain that $\{v,v_{1},v_{3}\}$ is complete to $V(S)$, and $(N(v)\setminus V(S))\cup\{v_{2},v_{4},v_{5}\}$ is anticomplete to $V(S)$.
+Suppose that $V(S)$ is not a homogeneous set of $G$.
+There exists a vertex $z$ which has a neighbor and a non-neighbor in $V(S)$.
+Clearly, $z\notin N(v)\cup\{v\}\cup V(C)$.
+Since $S$ is connected, it follows that there exists an edge $xy\in G[S]$ such that $z\sim x$ and $z\not\sim y$.
+To avoid an induced bull on $\{y,x,v_{1},z,v_{2}\}$, we have that $z\sim v_{2}$.
+Similarly, we have that $z$ is complete to $\{v_{2},v_{4},v_{5}\}$.
+So $z\in N(V(C))$.
+By Claim 4.1, $z$ has a neighbor in $\{v_{1},v_{3}\}$.
+By symmetry, we may assume that $z\sim v_{1}$.
+But then $\{v_{1},x,z,v,v_{4}\}$ induces a bull, a contradiction.
+Therefore, $V(S)$ is a homogeneous set of $G$.
+This completes the proof of Lemma 4.1.
+$\square$
+Now we begin to prove Theorem 1.3.
+Proof of Theorem 1.3: Since the graph $F$ is non-perfectly divisible, it suffices to prove that every (bull, $P_{6}$, $F$)-free graph are perfectly weight divisible.
+Suppose not, and let $G$ be a (bull, $P_{6},F$)-free MNWD graph.
+By Lemma 2.3, we have that $G$ is a connected odd net-free graph, and for every vertex $v\in V(G)$, $G[M(v)]$ is imperfect and contains no odd antihole of number of vertices at least 7.
+By Lemmas 2.5 and 2.6, we have that $\omega(G)\geq 3$ because $G$ is non-perfectly weight divisible.
+Let $v\in V(G)$ such that $v$ belongs to a maximum clique of $G$.
+By Theorem 1.1, we have that $G[M(v)]$ contains a 5-hole.
+Therefore, by Lemma 4.1, $G$ has a homogeneous set, which contradicts with Lemma 2.1.
+This completes the proof of Theorem 1.3.
+$\square$
+5 Proof of Theorems 1.5, 1.6 and 1.7
+In this section, we will prove Theorems 1.5-1.7.
 Substituting a vertex $v$ of a graph $G$ by a graph $H$ is an operation that creates a graph obtained from the disjoint union of $H$ and GvG-v by adding an edge between every vertex of $H$ and every neighbor of $v$ in $G$.
@@
 We denote by ${\cal C}_{s}$ the closure of $\cal C$} under taking substitution.
-We say a bull-free graph $G$ is basic if neither $G$ nor $\overline{G}$ contains an odd hole with both a vertex which is complete to the odd hole and a vertex which is anticomplete to the odd hole.
-Chudnovsky and Safra [8] proved the following two lemmas.
 Lemma 5.1
-((1.4 of [8] )) Every bull-free graph can be obtained via substitution from basic bull-free graphs.
+[…]
 Lemma 5.2
-((4.3 of [8] )) Let $G$ be a basic bull-free graph. Then for every vertex $v\in V(G)$, either $G[N(v)]$ is perfect or $G[M(v)]$ is perfect.
+[…]
 Lemma 5.3
-((6.1 of [2] )) Let $\cal C$} be a class of graphs.
+((Theorem 6.1 of [2] )) Let $\cal C$} be a class of graphs.
 If $\cal C$} is hereditary and polynomially $\chi$-bounded with $\chi\leq\omega^{k}$, then ${\cal C}_{s}$ is polynomially $\chi$-bounded with $\chi\leq\omega^{2k+3}$.
 Lemma 5.4
+((Lemma 5 of [4])) Let $G$ be a connected (bull, odd net)-free graph, and let $v\in V(G)$. Then $G[M(v)]$ contains no odd antihole of number of vertices at least 7.
+Lemma 5.5
 [17] The class of perfect graphs is closed under taking substitution.
-Lemma 5.5
-Let $G$ be a locally perfect $(P_{6},bull)$-free graph. Then there exists a partition $(X,Y)$ of $V(G)$ such that $G[X]$ and $G[Y]$ are perfect.
-Proof.
-We prove it by induction on $|V(G)|$.
+We now proceed to prove Theorems 1.5-1.7.
+Proof of Theorem 1.5: We prove the Theorem by induction on $|V(G)|$.
 It is certain that it holds when $|V(G)|$ is small.
@@
 Let $v\in V(G)$ such that $v$ belongs to a maximum clique of $G$.
-If $G[M(v)]$ is perfect, then $(N(v),M(v)\cup\{v\})$ is a partition of $V(G)$ which satisfies the Lemma since $G$ is locally perfect.
+If $G[M(v)]$ is perfect, then $(N(v),M(v)\cup\{v\})$ is a partition of $V(G)$ which satisfies the lemma since $G$ is odd net-free.
 Therefore, we may assume that $G[M(v)]$ is imperfect.
@@
 Recall that $1<|S|<|V(G)|$.
-Since $G$ is connected, we have that there exists a vertex which is complete to $S$, and so $G[S]$ is perfect because $G$ is locally perfect.
+Since $G$ is connected, we have that there exists a vertex which is complete to $S$, and so $G[S]$ is perfect as $G$ is odd net-free.
 Since $S\neq\mbox{{\rm\O }}$, we may assume that $x\in S$.
 Let $G^{\prime}=G[V(G)\setminus(S\setminus\{x\})]$.
-It is easy to verify that $G^{\prime}$ is a locally perfect $(P_{6}$, bull)-free graph and $|V(G^{\prime})|<|V(G)|$.
 Therefore, by inductive hypothesis, there exists a partition $(X^{\prime},Y^{\prime})$ of $V(G^{\prime})$ such that $G[X^{\prime}]$ and $G[Y^{\prime}]$ is perfect.
@@
 Therefore, the partition $(X,Y)$ satisfies the lemma.
-This completes the proof of Lemma 5.5.
+This completes the proof of Theorem 1.5.
 $\square$
-Lemma 5.6
-Let $G$ be a basic $(P_{6},bull)$-free graph. Then $\chi(G)\leq\omega(G)^{2}$.
-Proof.
-We prove it by induction on $|V(G)|$.
+Proof of Theorem 1.6: We prove the Theorem by induction on $|V(G)|$.
 It is certain that it holds when $|V(G)|$ is small.
@@
 We may assume that $\omega(G)\geq 2$ and $G$ is connected by inductive hypothesis.
-If $G$ is locally perfectly, then $\chi(G)\leq 2\omega(G)\leq\omega(G)^{2}$ by Lemma 5.5, and so we are done.
-Therefore, we may assume that $G$ is not locally perfect.
+If $G$ is odd net-free, then $\chi(G)\leq 2\omega(G)\leq\omega(G)^{2}$ by Theorem 1.5, and so we are done.
+Therefore, we may assume that $G$ contains an odd net.
 Now, by Lemma 5.2, there exists a vertex $v$ such that $G[M(v)]$ is perfect.
@@
 $\chi(G)\leq\chi(G[N(v)])+\chi(G[\{v\}\cup M(v)])\leq(\omega(G)-1)^{2}+\omega(G)\leq\omega(G)^{2}.$
-This completes the proof of Lemma 5.6. $\square$
-Now, we can prove Theorem 1.5.
-Proof of Theorem 1.5: Let $\cal G$} be the class of ($P_{6}$, bull)-free graphs, and $\cal C$} be the class of basic ($P_{6}$, bull)-free graphs.
-It is easy to verify that $\cal C$} is hereditary.
-By Lemmas 5.3 and 5.6, it suffices to show that ${\cal G}\subseteq{\cal C}_{s}$.
+This completes the proof of Theorem 1.6. $\square$
+Proof of Theorem 1.7: Let $\cal G$} be the class of (bull, $P_{6}$)-free graphs, and $\cal C$} be the class of (bull, $P_{6}$, odd net+)-free graphs.
+By Lemma 5.3 and Theorem 1.6, it suffices to show that ${\cal G}\subseteq{\cal C}_{s}$.
 Let $G\in{\cal G}$.
-Since $G$ is bull-free, by Lemma 5.1, we have that $G$ can be obtained from a basic bull-free graph $G_{1}$ by substituting a vertex by a basic bull-free graph $G_{2}$.
+Since $G$ is bull-free, by Lemma 5.1, we have that $G$ can be obtained from a (bull, odd net+)-free graph $G_{1}$ by substituting a vertex by (bull, odd net+)-free -free graph $G_{2}$.
 Since $G$ is $P_{6}$-free and $G_{1},G_{2}$ are induced subgraphs of $G$, we have that both $G_{1}$ and $G_{2}$ are $P_{6}$-free, which implies that $G_{1},G_{2}\in{\cal C}$.
@@
 This completes the proof of Theorem 1.7. $\square$
-Remark
-[…]
-More generally, let ${\cal G}_{1}$ be a hereditary class of graphs in which every trianlge-free graph is $k$-colorable, let ${\cal G}_{2}$ be the class of bull-free graphs, and let ${\cal G}={\cal G}_{1}\cap{\cal G}_{2}$.
-In [6], Chudnovsky et al.
-proved that $\cal G$} is polynomially $\chi$-bounded.
-Here, what are we more concerned about is that is it true that every graph in $\cal G$} is perfectly divisible if and only if $k=3$ ?
 Declarations